### v1

# Flexible-length Text Infilling for Discrete Diffusion Models

Abstract
Discrete diffusion models are a new class of text generators that offer advantages such as bidirectional context use, parallelizable generation, and flexible prompting compared to autoregressive models. However, a critical limitation of discrete diffusion models is their inability to perform flexible-length or flexible-position text infilling without access to ground-truth positional data. We introduce DDOT (Discrete Diffusion with Optimal Transport Position Coupling), the first discrete diffusion model to overcome this challenge. DDOT jointly denoises token values and token positions, employing a novel sample-level Optimal Transport (OT) coupling. This coupling preserves relative token ordering while dynamically adjusting the positions and length of infilled segments, a capability previously missing in text diffusion. Our method is orthogonal to existing discrete text diffusion methods and is compatible with various pretrained text denoisers. Extensive experiments on text infilling benchmarks such as One-Billion-Word and Yelp demonstrate that DDOT outperforms naive diffusion baselines. Furthermore, DDOT achieves performance on par with state-of-the-art non-autoregressive models and enables significant improvements in training efficiency and flexibility.
Flexible-length Text Infilling for Discrete Diffusion Models
Andrew Zhang, Anushka Sivakumar, Chiawei Tang, Chris Thomas Virginia Tech, Virginia, USA azhang42vt.edu
1 Introduction
While autoregressive models are effective, they also involve highly sequential sampling, cannot take bidirectional context, and lead to constrained architectures by requiring a decoder mask. In contrast, discrete diffusion models Lou et al. ; Sahoo et al. (2024) can parallelize generation by denoising multiple tokens simultaneously, use bidirectional information, do not need a decoder mask, and, notably, allow for more controllable generation. Previous work Gat et al. ; Shi et al. (2024) has demonstrated that discrete diffusion models are capable of arbitrary prompting locations, whereas autoregressive models are only capable of left-to-right text completion. However, this advantage has a significant limitation: existing diffusion models cannot alter the distances between these prompt tokens. Consequently, existing text diffusion models cannot generate the ground-truth sample without access to the oracle positions of the prompt and infilled text.
Figure 1: Our novel diffusion across token positions enables dynamic token movement for infilling. Unlike prior methods, DDOT learns to move mask tokens to appropriate locations, such as to the right of "brown," even if initially unmasked there. The OT coupling (colored lines) significantly simplifies this learning by drastically reducing possible permutations.
We solve this issue by enabling discrete diffusion models to learn where to move tokens. Specifically, we design a diffusion process that operates across token positions, allowing the model to vary the position and length of infilled spans. Furthermore, given the importance of token positioning in preserving semantic meaning He et al. , we incorporate sample-level OT (optimal transport) coupling to maintain relative token ordering throughout the diffusion process. Even minor positional changes can dramatically alter meaning, as seen in phrases like "The child’s green coat" and "The green child’s coat". DDOT’s OT coupling preserves this relative ordering throughout the diffusion process while supporting flexible-length text infilling. Our OT coupling prevents such swaps and drastically improves DDOT’s downstream performance across all studied benchmarks and metrics. Extensive experiments show that DDOT outperforms naive diffusion baselines and achieves on par performance with state-of-the-art non-autoregressive (NAR) models.
In summary, our contributions are as follows:
  • We propose DDOT, the first discrete text diffusion method for infilling arbitrary text sequences without ground-truth span lengths.
  • We provide extensive experiments on DDOT that show it achieves outperforms diffusion baselines and achieves performance on par with state-of-the-art NAR models.
  • We provide detailed ablations and visualizations that verify DDOT’s effectiveness for adjusting the position and length of infilled text spans and provide insights into our novel sample-level OT coupling. The OT coupling significantly outperforms naive diffusion across all tested benchmarks and metrics.
Figure 2: DDOT learns to vary infilled span lengths and positions unlike prior fixed-position diffusion methods. (Left) We compute two separate intra-set OT couplings within the prompt positions and the response positions. This constraint drastically simplifies possible permutations. Right Given a timestep $t$, we predict the token and position.
2 Related Work
2.1 Lexicographically Constrained Generation
Constrained text generation has been explored through a variety of approaches, including AR and NAR methods Zhang et al. (2020); Iso (2024); He (2021); Stern et al. (2019); Lu et al. (2022). POINTER Zhang et al. (2020) enables flexible token generation through iterative insertion, though it still depends on sequential token prediction and multiple forward passes. AutoTemplate Iso (2024) approaches constrained text generation by simply feeding the prompt tokens into an encoder-decoder style model. CBART He (2021) extends the POINTER architecture by moving to an encoder-decoder model. Autoregressive methods, while effective for their specific use cases, inherit fundamental limitations: they require sequential generation that scales linearly with sequence length, and their causal attention masks prevent full utilization of bidirectional context during generation. Most critically, for text infilling tasks, these approaches struggle to simultaneously consider both past and future context when generating intermediate content Cao et al. (2023).
2.2 Discrete Diffusion Models
Discrete diffusion models offer an innovative approach to text generation, addressing key limitations of autoregressive methods Lou et al. ; Ren et al. (2024); Gong et al. (2024); Sahoo et al. (2024). These models denoise corrupted text sequences, enabling parallel updates of multiple tokens rather than the token-by-token process of autoregressive methods, reducing the number of forward passes. Additionally, their bidirectional nature allows them to leverage both past and future context in contrast to causal masking that constrains autoregressive models. Early frameworks like D3PM Austin et al. adapted continuous diffusion to discrete tokens using Markovian corruption, while subsequent advances such as multinomial diffusion Hoogeboom et al. (2021) improved noise schedules and training efficiency.
Recent work on score-based discrete diffusion has further advanced the field by providing analytical solutions for the denoising process. Instead of directly modeling transition probabilities SEDD Lou et al. uses a score-based approach that learn the gradient of the log probability, which proves particularly effective for handling high-dimensional categorical data like text.
However, despite these advantages, current discrete diffusion models face a significant limitation: they require fixed token positions throughout the generation process. This constraint makes them unsuitable for flexible-length text infilling, where the length of the generated text might differ from the original masked region.
2.3 Optimal Transport Coupling for Flexible Generation
OT Villani et al. (2009) coupling has been well studied in image generation through theoretical foundations such as Rectified Flow, Liu et al. , which showed how continuous normalizing flows can be understood as OT paths. This framework demonstrated that probability mass can be efficiently transported through straight trajectories while preserving structural relationships, providing insights into how sequential data can be manipulated while maintaining order.
Building on these insights, Tong et al. showed how OT can be incorporated into diffusion models to accelerate generation. By defining appropriate transport costs between positions, OT coupling ensures that dimensions within a sample move in straight, predictable paths.
Although previous works incorporated OT into diffusion models with a focus on generation speed, our work leverages these properties to enable flexible-length text infilling. By coupling the discrete token values with continuous position variables through OT, we allow the model to optimize both token content and positioning simultaneously. This approach maintains the parallel generation and bidirectional context advantages of discrete diffusion while adding the ability to dynamically adjust sequence lengths. The OT coupling effectively serves as a bridge between the discrete token space and continuous position space, enabling natural language generation that respects both local token relationships and global sequence structure.
3 Background: Masked Text Diffusion
Discrete diffusion adapts the continuous nature of diffusion processes to text. The simplest and most performant version of discrete diffusion is masked diffusion. Rather than adding gaussian noise to continuous values such as pixels, masked text diffusion assigns a certain probability of masking tokens throughout the forward diffusion process. For the sake of this paper, masked diffusion can be seen as a masked language model (like BERT Devlin et al. (2019)) that works at gradually decreasing masking ratios to generate text. Specifically, our text diffusion process follows Score Entropy Discrete Diffusion (SEDD) Lou et al. , modeling a score function across a support of $N$ states or token values. The forward diffusion process is given by a continuous time Markov chain (CTMC):
$\frac{dp_{t}}{dt}=Q_{t}p_{t}\quad p_{0}\approx p_{\mathrm{data}}$ (1)
$Q_{t}\in\mathbb{R}^{n\times n}$ denotes a transition matrix where the columns give the probability of transitioning from one discrete state (or token) to another. In masking diffusion models, the diagonal is $-1$, the last row is $1$, and all other elements are 00, moving probability mass from any token value to the last token in the vocabulary, the mask token. To extend the process to sequences rather than individual tokens, SEDD applies discrete diffusion independently to all tokens in a sequence.
SEDD reverses the textual diffusion equation by modeling the score $s_{\theta}(x)_{y}$ of transitioning from token $x$ to token $y$ with a score entropy loss:
${\scriptstyle\mathcal{L}_{\text{tok}}=\mathbb{E}_{(x,x_{0})}\Big{[}\sum_{y\neq x}w_{xy}\Big{(}s_{\theta}(x)_{y}-\frac{p(y|x_{0})}{p(x|x_{0})}\log s_{\theta}(x)_{y}\Big{)}\Big{]}}$ (2)
where $w_{xy}$ weighs the importance of different states. Finally, to simulate the reverse diffusion process, we either take Euler steps or use an analytical denoiser Lou et al. .
4 Approach
Previous discrete diffusion models must know the ground truth positions of mask tokens to infill a sequence, as the positions of tokens are fixed throughout the entire diffusion process. To address this limitation, we introduce DDOT, a diffusion framework that jointly denoises discrete token values and continuous token positions, allowing for flexible infilling spans.
4.1 Continuous Position Diffusion
DDOT gradually transforms token positions from a simple initial distribution to their ground truth permutation. To achieve this, we denoise $l$ token positions $z_{t}\in[-1,1]^{l}$ in the continuous domain. The initial (limiting) distribution of positions is sampled from a uniform distribution, $z_{T}\sim\mathcal{U}(-1,1)^{L}$, where $L$ represents the maximum sequence length. The ground truth positions are defined as:
$z_{0}=linspace\left(-\frac{l}{L},\frac{l}{L},l\right),$ (3)
where $z_{0}$ is evenly spaced and scaled to match the length of the true sequence. This setup ensures that the position diffusion process captures the gradual transition from a simple prior to a structured output aligned with the token ordering and length. Similar to fixed position models, we scale by $l$ to provide information on the abosolute distances between tokens.
Our preliminary studies reveal that naively performing diffusion on the token positions leads to poor performance. We observe that adding noise to token positions destroys the relative ordering of text and adds a combinatorially exploding number of permutations, problems not present in absolute-valued continuous domains such as image generation Ho et al. (2020); Karras et al. (2022). Since the relative ordering of text can drastically change its meaning and consequently abruptly destroy the signal across timesteps, it is crucial to fix token ordering across time.
Unlike traditional diffusion processes that aim to completely destroy the data signal at the final timestep $T$, we provide the ordering of the prompt tokens at $T$ to the model. To maintain the prompt token ordering, we introduce a sample level OT coupling between the initial $z_{0}$ and final $z_{T}$ token positions of the diffusion process. Additionally, we do not add noise to the positions and instead use linear interpolation between $z_{0}$ and $z_{T}$.
Following Liu et al. ; Albergo and Vanden-Eijnden , we model the forward position diffusion process with an ordinary differential equation (ODE)
$dZ_{t}=v(Z_{t},t)dt$ (4)
where $v(z_{t},t)$ is the velocity (rate of change) of the positions at time $t$. We partition $x$ and $z$ into prompt ($x^{p}$, $z^{p}$) and infilled response ($x^{r}$, $z^{r}$) subsets. The limiting positions $z_{T}$ are determined as the solution to the linear sum assignment problem, minimizing OT cost after scaling $z_{0}$in $[-1,1]$:
$z^{*}_{T}=\pi(\frac{z_{0}L}{l},z_{T}),\quad\pi=\arg\min_{\pi^{\prime}\in S_{n}}\sum_{i=1}^{n}C_{i,\pi^{\prime}(i)}$ (5)
where $C_{i,j}$ is the Euclidean distance between positions and $S_{n}$ is all possible permutations of length $n$. For prompt tokens, the optimal transport is balanced, so $|z^{p}_{T}|=|z^{p}_{0}|=l^{p}$ However, the number of infilled tokens in the limiting distribution is almost always larger than that of the ground truth $|z^{r}_{T}|=L-l^{p}\geq l-l^{p}=|z^{r}_{0}|$ To reconcile the different sizes of $z_{0}$ and $z_{T}$, we treat the remaining unassigned $L-l$ positions as pad tokens. Specifically, we set the ground truth tokens to the pad token and the ground truth positions to stationary paths before scaling: $z^{pad}_{0}=z^{pad}_{T}*\frac{l}{L}$. Importantly, these stationary paths maintain that the coupling between $z_{0}$ and $z_{T}$ achieves the optimal OT cost.
This approach establishes a coupling within prompt and infilled tokens, ensuring that tokens in each subset maintain non-crossing paths, a key property of OT Tong et al. ; Villani et al. (2009). While intra-set paths remain ordered to minimize transport cost, DDOT allows inter-set crossings (i.e., between prompt and infilled tokens), enabling flexible token positioning while preserving the relative order of prompt tokens. Path visualizations are available in A.4.
Beyond prompt tokens, DDOT ensures that relative token order within all sets, prompt and infilled response, aligns with their ground truth at any timestep. This guarantees smooth transitions and avoids disruptions caused by abrupt positional swaps. Using OT, DDOT simplifies position manipulation, reducing unnecessary degrees of freedom.
OT is traditionally computationally expensive, requiring iterative calculations on the CPU. However, our OT coupling is computationally efficient, addressing the large-scale data demands of generative text modeling. Previous methods approximate dataset-level OT through per-batch computations Tong et al. , typically relying on algorithms like the Sinkhorn method Cuturi (2013) and operating in high-dimensional spaces (e.g., image latents). In contrast, our OT coupling operates at the scalar level and limited to $L$ elements, corresponding to the model’s context size—a scale significantly smaller than the number of samples in a dataset. For prompt tokens, the OT coupling is balanced and can therefore be efficiently computed by simply sorting $z_{0}$ and $z_{T}$. This efficiency enables us to calculate exact sample-level OT, unlike the approximations required in previous dataset-level approaches.
For training, we randomly sample a timestep $t$ and retrieve the corresponding positions in a simulation-free manner with linear interpolation $z_{t}=(1-t)z_{0}+tz_{T}$. The training objective is a weighted mean-squared error loss:
${\mathcal{L}_{pos}(\theta)=\mathbb{E}_{(z,t)}\left[Q_{t}(x_{t},y)\|v_{\theta}(z_{t},t)-(z_{0}-z_{T})\|^{2}\right]}$ (6)
We investigate two methods of initializing $z_{T}$ to begin generation. The first method is to sample $z_{T}\sim\mathcal{U}(-1,1)^{L}$, which we refer to as DDOT-random (DDOT-R). We then randomly select $l^{p}$ tokens from $z_{T}$ to serve as the positions for the prompt tokens. However, some areas of high density arise when randomly sampling which tend to map to pad tokens because there are less tokens in the corresponding ground truth region. Therefore, we propose DDOT-uniform (DDOT-U), which uniformly spaces out the prompt and infilled positions: $z^{p}_{T}=\text{linspace}(-1,1,l^{p})$ and $z^{r}_{T}=\text{linspace}(-1,1,l^{r})$.
During the sampling process, the token paths tend to follow straight trajectories because the ground truth velocity, $z_{0}-z_{T}$, remains constant over time. Therefore, we find it sufficient to use Euler steps during sampling to predict the positions. Combined with $\tau$-leaping Lou et al. , the straight position paths allow for favorable compute-accuracy tradeoffs and result in fast sampling. The total loss $\mathcal{L}$ is a simple linear combination $\mathcal{L}=\mathcal{L}_{tok}+\lambda\mathcal{L}_{pos}$.
4.2 Simultaneous Text & Position Diffusion
DDOT performs discrete text and continuous position diffusion simultaneously, as these processes operate independently in continuous time. We therefore predict both token value scores and position velocities in a single forward pass. This independence also enables simulation-free training by independently sampling token and position states at arbitrary timesteps. We summarize the training procedure in Appendix A.5
4.3 Implementation Details
Due to compute constraints, we implement our framework on SEDD Lou et al. . Our flexible position diffusion component is orthogonal to existing text diffusion methods and minimally alters the permutation of text tokens for infilling. This design allows DDOT to integrate seamlessly with various pretrained text diffusion models.
We extend SEDD, which is based on the Diffusion Transformer architecture Peebles and Xie (2023), with two additional modules. First, we introduce a learnable type embedding applied directly after the token embedding lookup. This embedding indicates whether a token is part of the prompt or the masked response ($x\in x^{p}$ or $x\in x^{r}$), which is critical for assigning each token to the correct OT flow. Second, we add a linear head at the end of the diffusion transformer to compute $v_{\theta}(z_{t},t)$.
To incorporate continuous positional information, we scale $z_{t}$ from the range $[-1,1]$ to match the context length of the original pretrained model (1024). We then use Rotary Position Embeddings Su et al. (2024), a standard technique in discrete diffusion models. Implementation details can be found in subsection A.3.
5 Experiments
5.1 Experimental Setup
Datasets    We evaluate our approach on the One-Billion-Words and Yelp, following the preprocessing steps outlined in prior works on infilling and lexicographically constrained generation Miao et al. (2019); Zhang et al. (2020); Iso (2024). These datasets consist of examples with $1$ to $6$ keywords that must be infilled while maintaining their relative order to generate coherent sentences. In addition to randomly masking positions, we also introduce the "block" masking method that masks a single continuous chunk of text 00 to $L/2$ tokens long (32 for one-billion-word and yelp and 512 for CodeParrot). Finally, we apply the aforementioned masking methods to the Python subset of the CodeParrot dataset. Table 1 illustrates examples of this lexicographically constrained generation task.
Keywords: earned, cents, share, costs
One-Billion-Word-Random: It earned 28 cents a share , excluding restructuring costs.
Keywords: we are so incredibly wedding .
Yelp-Block: we are so incredibly happy we chose this venue for our wedding .
Table 1: Example generations for the keywords-to-sentence generation on One-Billion-Word and Yelp.
Model One-Billion-Word-Block Yelp-Block CodeParrot-Block CodeParrot-Random
B2/B4 N2/N4 M SR B2/B4 N2/N4 M SR B2/B4 CB SR B2/B4 CB SR
LC 55.1/47.9 4.38/4.40 37.4 54.7 59.1/51.7 7.77/8.48 37.9 60.3 51.0/49.4 53.3 16.9 15.4/24.8 19.4 0.
PoP 59.2/50.0 8.43/8.77 37.1 58.6 64.9/48.6 7.81/8.47 38.1 67.1 29.4/13.3 21.7 0.34 12.5/2.75 13.9 0.
DDOT-R 62.2/57.1 7.42/8.16 42.1 99.7 62.5/57.3 7.50/8.24 42.2 99.8 47.2/41.1 38.5 1.01 57.2/43.5 39.2 3.53
DDOT-U 63.7/58.4 8.40/8.79 42.1 100. 64.9/59.5 8.06/8.86 42.9 100. 53.7/51.0 50.5 10.8 58.2/40.4 45.4 11.2
Table 2: DDOT outperforms diffusion baselines on standard sequences (0-32 prompt tokens). Metrics are BLEU (B2, B4), NIST (N2, N4), METEOR (M), and Success Rate (SR). Top scores are bolded.
Model One-Billion-Word-Random Yelp-Random
B2/B4 N2/N4 M SR B2/B4 N2/N4 M SR
Autoregressive Models
GBS Hokamp and Liu (2017) 10.1/2.8 1.49/1.50 13.5 $\leq 100.$ 13.6/4.5 1.68/1.71 15.3 $\leq 100.$
InstructGPT Ouyang et al. (2022) 10.1/2.8 1.72/1.73 13.0 92.33 9.3/2.4 1.42/1.44 13.6 92.17
AutoTemplate-small Iso (2024) 16.4/6.1 3.11/3.15 15.5 100. 22.5/9.5 3.51/3.63 17.1 100..
AutoTemplate-base Iso (2024) 18.3/7.6 3.39/3.45 16.0 100. 23.7/10.8 3.62/3.76 17.8 100..
AutoTemplate-large Iso (2024) 18.9/8.1 3.49/3.54 16.2 100. 24.1/11.1 3.68/3.83 17.9 100..
Non-Autoregressive Models
Traditional Models
   SeqBF Mou et al. (2016) 4.4/0.7 0.62/0.62 7.0 $\leq 100.$ 6.9/2.1 0.52/0.53 8.7 $<100.$
   CGMH Miao et al. (2019) 9.9/3.5 1.15/1.17 13.1 100. 12.3/4.6 1.41/1.45 14.6 $\leq 100.$
   POINTER Zhang et al. (2020) 8.7/1.6 2.11/2.12 14.3 100. 10.6/2.4 2.14/2.16 16.8 100.
   CBART He (2021) 15.6/6.6 2.16/2.19 15.2 100. 19.4/9.0 2.54/2.64 17.4 100.
Diffusion Models
   LC 15.36/6.52 2.02/2.05 14.99 99.82 20.55/9.66 2.76/2.87 17.58 99.75
   PoP 16.59/5.66 3.06/3.09 15.15 99.05 21.13/7.97 3.31/3.40 17.90 98.87
   DDOT-R (Ours) 15.7/5.1 3.17/3.21 14.6 99.60 19.7/7.0 3.26/3.34 17.4 99.77
   DDOT-U (Ours) 16.3/5.2 3.13/3.17 15.0 100. 21.2/7.9 3.43/3.52 17.7 100.
Table 3: DDOT performs on-par with state-of-the-art NAR models on short sequences (1–6 prompt tokens). Top NAR scores are bold; second-best are underlined. Since SOTA NAR backbones (e.g. diffusion) still lag behind AR backbones, we focus on NAR comparisons.
Figure 3: Success rate on block datasets. LC and PoP increasingly generate invalid responses (missing or swapping prompt tokens) as the number of prompt tokens grows.
Training Details    To align the position prediction modules, we first finetune SEDD with the added modules on FineWeb-Edu Penedo et al. (2024). Afterwards, we further finetune on One-Billion-Word and Yelp datasests. For simplicity, we always keep all parameters unfrozen and simultaneously optimize both $\mathcal{L}_{tok}$ and $\mathcal{L}_{pos}$.
In line with SEDD, we train our model in two configurations: small (90M non-embedding parameters) and medium (320M non-embedding parameters). DDOT-medium is on the same scale as CBART (406M parameters), and AutoTempalate-base (220M parameters). Following SEDD, we use the AdamW optimizer with a learning rate of 3e-5. We set $\lambda=10$ when using scaling. For each experiment, we either use 48 L40s (48 GB) GPUs, 80 A30 (24GB) GPUs, or 8 A100 (80GB) GPUs.
Baselines    We compare our method against strong autoregressive (AR) and non-autoregressive (NAR) baselines. AutoTemplate Iso (2024), the state-of-the-art AR model, leverages the T5 Raffel et al. (2020) family of pretrained models. Specifically, AutoTemplate parses the lexicographically constrained generation task into a template that is autoregressively generated from left to right. The previous state-of-the-art NAR method, CBART He (2021), is built upon the BART Lewis et al. (2020) pretrained framework and iteratively inserts tokens into a sequence.
We also introduce two diffusion-based models that follow the same training procedure as DDOT. Left Context (LC) concatenates all the prompt tokens to the left of the sequence and generates the response to the right of a separator token. Position Prediction (PoP) uses a SEDD model with a linear head that first predicts the positions of every token. Then, this sequence is fed through a finetuned fixed-position SEDD.
Distribution Annealing    Many lexiocagphically constrained generation baselines including AutoTemplate and CBART use distribution annealing methods such as top-p, top-k, greedy sampling, and beam search. To provide a parallel to greedy decoding which always takes the top token probability, we anneal the distributions of our tokens values during sampling to only include the most probable token. Specifically, given the predicted probability of a certain token being mask $\hat{p}(x^{mask})$, we assign $1-\hat{p}(x^{mask})$ to the token value with the highest probability excluding the mask token. The rest of the token probabilities are set to 00. Greedy decoding in prior models (such as autoregressive) is deterministic, collapsing the tree of all generation paths into a single path. However, our annealing process maintains generation diversity (A.2) because the model must still sample from the annealed distribution with the top token value and the mask token. Whenever possible, we evaluate against the greedy decoding baseline.
Metrics  Following prior works Miao et al. (2019); Zhang et al. (2020); Iso (2024), we evaluate on BLEU-2/4 Papineni et al. (2002), NIST-2/4 Doddington (2002), and METEOR-v1.5 Denkowski and Lavie (2014), and success rate (SR).
5.2 Main Results
We present lexicographically constrained generation results in Table 2. Our approach uses greedy annealing and is compared against greedy decoding wherever applicable, including the CBART greedy decoding baseline. Our method achieves competitive performance with previous NAR models, approaching AR performance. Notably, our model achieves state of the art performance on most metrics among the diffusion baselines. Our method does well on block infilling, which may be more useful in real-world applications. Furthermore, we notice that DDOT scales well to longer sequences. Specifically, it frequently generates valid responses that include all the prompt words in the same relative ordering, as shown in the success rates (SR). In contrast, diffusion baselines quickly generate invalid responses as the number of prompt tokens increase (Figure 3, Table 2). However, in benchmarks with 6 or less prompt tokens, diffusion baselines maintain high SR (Table 3). This may be because the fixed-position models have room to correct generation when the ratio of prompt to response tokens is low.
Table 3 Compares results with previous works in the lexicographically constrained generation line of work. Since pretrained diffusion models lag behind AR models, an issue not unique to DDOT, we elect to focus on NAR models. DDOT performs on par with the previous SOTA models, and achieves higher SR than all diffusion baselines. Although DDOT underperforms LC and PoP in some metrics, we argue that one-billion-word-random and yelp-random over-index on the unrealistic task of generating text with only $1-6$ randomly spaced tokens, and that Figure 3 shows the broader trend when DDOT is scaled to more prompt tokens.
5.3 Analysis
In this section we investigate the effect of random versus uniform position initialization, the inclusion of OT coupling, and the impact of varying the number of sampling steps.
Figure 4: (a) and (b) exhibit Ground Truth Token Velocities. The velocities without OT (a) have a lot of crossing lines demonstrating instability in matching whereas the velocities with OT coupling (b) show almost straighter lines throughout the denoising process. (c) Performance tends to increase with sampling steps.
Position Initialization    In Table 2 and Table 3 we also explore the difference between DDOT-R and DDOT-U. In DDOT-R, pad tokens tend to cluster in areas of high density because the OT finds no match for them. However, the pad tokens in DDOT-U tend to be evenly spaced out. We find that DDOT-U consistently outperforms DDOT-R.
OT Coupling    To demonstrate the importance of OT coupling between source and target positions, we retrain the small version of DDOT without OT coupling and provide a quantitative comparison in Table 4. Models trained with OT coupling consistently outperform those using independent (random) coupling. We theorize that the OT coupling provides more signal about the ordering of tokens throughout the diffusion process. Specifically, the DDOT guarantees that the relative ordering of the prompt and generated tokens at any timestep is the same as the original ordering. In contrast, independent coupling requires the model to infer the original ordering of tokens—a challenging task given the numerous plausible orderings that can result from interspersed prompt tokens.
OT? B2 B4 N2 N4 M
One-Billion-Word
No OT 13.2 4.6 1.60 1.62 14.1
OT 15.7 5.1 3.2 3.2 14.6
Yelp
No OT 16.39 5.94 2.05 2.10 16.0
OT 18.6 6.9 3.13 3.21 16.6
Table 4: Our OT coupling drastically improves preformance across all metrics. Ablation on OT coupling with small model size.
Position Over Time    We qualitatively compare the ground truth token paths during training Figure 4. With OT coupling, token trajectories exhibit significantly fewer crossings, maintaining relative order throughout the generation process. In contrast, the independent coupling frequently permutes tokens. Visualizations of token paths during inference are available in subsection A.4
Number of Sampling Steps    One advantage of diffusion models over autoregressive models is their ability to exchange compute for accuracy by varying the number of inference steps. Figure 4 shows how the number of sampling steps influences lexicographically constrained generation performance. As the number of sampling steps increases, performance also increases.
Wall Time Analysis    We evaluate the inference speed of DDOT against the diffusion baselines on the One-Billion-Word dataset. Table 5 presents the wall-clock time per batch alongside BLEU-2 and BLEU-4 scores for an increasing number of sampling steps.
DDOT demonstrates significantly better efficiency. For any given number of sampling steps, DDOT is not only faster than LC and competitive with PoP in terms of raw speed, but also achieves substantially higher BLEU scores. Notably, LC must regenerate prompt tokens and therefore requires up to double the input sequence length. PoP also requires an additional forward pass to predict initial positions.
Efficiency Considerations
Model Number of Sampling Steps
2 4 8 16 32 64
LC (Left Context)
Time (s/batch) 0.71 1.43 2.86 5.72 11.5 22.9
BLEU-2 38.3 44.3 50.7 53.3 54.5 55.1
BLEU-4 27.5 33.9 41.9 45.3 47.0 47.9
PoP (Position Prediction)
Time (s/batch) 0.53 0.90 1.64 3.13 6.09 12.0
BLEU-2 45.4 59.5 59.5 59.4 59.3 59.2
BLEU-4 40.7 50.3 50.3 50.2 50.1 50.0
DDOT-Uniform (Ours)
Time (s/batch) 0.44 0.81 1.54 3.03 5.98 12.0
BLEU-2 59.4 61.2 62.4 63.1 63.5 63.7
BLEU-4 55.5 56.8 57.5 58.0 58.2 58.4
Table 5: DDOT achieves superior BLEU scores with faster inference times. Inference speed (seconds per batch) and BLEU scores on One-Billion-Word for varying numbers of sampling steps.
The added modules to enable position prediction are lightweight, consisting of a linear head and two type embeddings. On the other hand, the LC baseline requires double the context length of DDOT because it must regenerate prompt tokens.
The OT calculation is highly efficient, taking 16 minutes and 11 seconds on an Intel Xeon 8462Y+ 64 core processor for the 10 billion token subset of FineWeb EDU. In practice, we stream the dataset, caching several OT couplings in advance without needing to preprocess the OT. With caching, it takes 4 minutes and 30 seconds to do 1000 training steps on an L40 gpu with a batch size of 256. Without caching, it takes 4 minutes and 27 seconds, a negligible difference.
6 Conclusion
In this work we introduce DDOT, the first discrete diffusion model capable of flexible-length text infilling by jointly denoising token values and positions. By incorporating optimal transport coupling, DDOT preserves token order while enabling dynamic position adjustment, addressing limitations in existing text diffusion models. Our experiments show that DDOT outperforms diffusion baselines and matches SOTA non-autoregressive models.
7 Limitations
DDOT inherits many drawbacks from previous discrete text diffusion methods Lou et al. ; Austin et al. ; Sahoo et al. (2024). First, diffusion backbones still lag behind autoregressive backbones. Since diffusion for text is an emerging field, this performance gap is expected. Similarly, the finetuned versions of these models (such as DDOT vs AutoTemplate) also exibit the same performance gap. Second, unlike autoregressive models, existing diffusion models cannot use a kv-cache to store previous activations. DDOT also inherits risks common between large-scale text models.
8 Ethics Statement
Artifacts  The artifacts that we used had public use licenses. Furthermore, we plan to release our code artifacts for public use after acceptance.
Dataset Considerations  We use publicly available datasets that underwent safety checks such as FineWeb-EDU.
Documentaion of Artifacts  Our work generates English text, and our codebase is primarily in Python.
Use of AI Assistants  We used AI assistants to help write our code and revise our paper.
Package Usage  we use nltk for BLEU, NIST, METEOR, and N gram diversity. We use https://pypi.org/project/codebleu/ for codebleu.
Appendix A Appendix
This appendix provides supplementary material to support the findings and methodology presented in the main paper. It includes detailed model size comparisons, generation diversity analysis, hyperparameter sensitivity studies, visualizations of token path dynamics, and the DDOT training algorithm.
A.1 Size Comparison
Table 6 presents a detailed breakdown of performance metrics (BLEU-2/4, NIST-2/4, and METEOR) for different sizes of our DDOT model (small and medium) and its variants (DDOT-R and DDOT-U). These results are shown for the One-Billion-Word-Random and Yelp-Random benchmarks. This comparison allows for an assessment of model scalability and the impact of the $z_{T}$ initialization strategy (random vs. uniform) across different model capacities. The medium-sized DDOT-U variant generally demonstrates the strongest performance, highlighting the benefits of uniform initialization and increased model parameters.
Model One-Billion-Word-Random Yelp-Random
B2 B4 N2 N4 M B2 B4 N2 N4 M
DDOT-R small (Ours) 15.7 5.1 3.16 3.19 14.6 18.6 6.9 3.13 3.21 16.6
DDOT-U small (Ours) 15.5 4.9 2.94 2.97 15.0 18.9 7.1 2.99 3.07 17.0
DDOT-R medium (Ours) 15.7 5.1 3.17 3.21 14.6 19.7 7.0 3.26 3.34 17.4
DDOT-U medium (Ours) 16.3 5.2 3.13 3.17 15.0 21.2 7.9 3.43 3.52 17.7
Table 6: Results on One-Billion-Word and Yelp. Metrics are BLEU (B2, B4), NIST (N2, N4), and METEOR (M). Best results are highlighted in bold.
A.2 Generation Diversity
Table 7 provides an evaluation of generation diversity for DDOT and baseline models on the One-Billion-Word-Random and Yelp-Random datasets. Diversity is measured using D2 and D4 metrics, which represent the percentage of unique bi-grams and 4-grams in the generated outputs, respectively. These metrics are crucial for assessing the variety and richness of the generated text beyond simple n-gram overlap with references. The results show how DDOT variants compare to other baselines like CBART, LC, and PoP in this regard, with CBART exhibiting high diversity scores. DDOT models offer competitive diversity, particularly DDOT-U on the One-Billion-Word dataset.
Model One-Billion-Word-Random Yelp-Random
D2 D4 D2 D4
CBART He (2021) 49.2 82.1 38.1 92.6
LC 37.5 89.4 21.5 69.0
PoP 38.85 91.4 19.9 68.3
DDOT-R (Ours) 30.7 85.3 20.2 67.7
DDOT-U (Ours) 34.03 88.2 18.6 66.6
Table 7: Results on One-Billion-Word and Yelp dataset for generation diversity. Metrics include Diversity (D2, D4). Best results are highlighted in bold.
A.3 Hyperparameter Sensitivity Analysis
This section details the hyperparameters explored and used in our experiments, providing transparency and aiding reproducibility.
A.3.1 $\lambda$ values
The choice of the loss weight $\lambda$ is critical as it balances the contribution of the token prediction loss ($\mathcal{L}_{\text{tok}}$) and the position prediction loss ($\mathcal{L}_{\text{pos}}$). Table 8 investigates the impact of different $\lambda$ values on model performance metrics (B2, B4, N2, N4, M) for the One-Billion-Word dataset. The results demonstrate the model’s performance across a range of $\lambda$ values, indicating a degree of robustness. For the experiments reported in this table, a $\lambda$ value around 10 appears to offer a good balance.
$\lambda$ B2 B4 N2 N4 M
3 63.07 58.27 7.56 8.32 42.64
10 63.04 58.23 7.55 8.30 42.85
30 63.64 58.63 7.73 8.50 42.66
100 63.05 58.18 7.56 8.32 42.6
300 63.25 58.24 7.63 8.39 42.50
1000 63.53 58.46 7.74 8.51 42.50
3000 62.95 57.95 7.59 8.34 42.27
Table 8: Impact of different $\lambda$ values on model performance metrics.
A.3.2 Training Hyperparameters
Table 9 lists the primary hyperparameters used for training the DDOT models presented in the main paper. These settings, including batch size variations (dependent on GPU resources like A30, L40, or A100), the $\lambda$ value used with scaling (300), weight decay, learning rate, optimizer choice (AdamW), and its $\beta$ parameters, are provided for reproducibility. These parameters were selected based on preliminary experiments and common practices in training diffusion models. Most parameters were selected based on SEDD’s values. $\lambda$ was emperically determined from small test runs.
Hyperparameter Value
Batch size 128 (A30) / 256 (L40) / 512 A(100)
$\lambda$ 300 (with scaling)
Weight decay 0.0
Learning rate $43e-$ 4
Optimizer AdamW
$\beta_{1}$ 0.9
$\beta_{2}$ 0.999
Table 9: Training hyperparameter specifications.
A.4 Path Visualizations
The following figures provide qualitative insights into the token path dynamics during the diffusion process. These visualizations help to understand the behavior of token positions during inference (sampling) and the nature of the ground truth paths the model learns during training, both with and without Optimal Transport (OT) coupling.
Figure 6 visualizes token trajectories during the inference (sampling) process without OT coupling, shown for both scaled (a) and unscaled (b) positions across a number of sampling steps. These plots illustrate how token paths can become less orderly and may curve or cross significantly towards the end of the generation process when OT is not employed to maintain relative order, potentially leading to permutation errors.
(a) Scaled positions over time without OT
(b) Unscaled positions over time without OT
Figure 5: Ablations on number of sampling steps. Token paths tend to curve towards the end of generation without OT coupling.
Figure 6, in contrast, shows token trajectories during inference (sampling) *with* OT coupling, again for scaled (a) and unscaled (b) positions. These paths are visibly straighter and more ordered throughout the generation process. This demonstrates the effectiveness of OT in maintaining relative token order and guiding tokens towards their target positions more directly, leading to more stable and accurate generation.
(a) Scaled positions over time with OT
(b) Unscaled positions over time with OT
Figure 6: Ablations on number of sampling steps. Token paths tend to be straighter towards the end of generation with OT coupling.
Figure 7 depicts the ground truth token paths between $z_{T}$ (initial noisy state) and $z_{0}$ (target state) during training, scaled between -1 and 1. Subfigure (a) shows these paths *without* OT coupling; here, $z_{T}$ represents a more arbitrary mapping from initial random positions to $z_{0}$, leading to many crossing paths. This illustrates the complex, potentially tangled mapping the model would need to learn. Subfigure (b) shows the paths *with* our sample-level OT coupling. In this case, $z_{T}$ is aligned with $z_{0}$ via optimal transport (within prompt and response sets), resulting in significantly straighter, non-crossing (within-set) trajectories. This comparison highlights how OT simplifies the learning objective for the positional diffusion component by providing more direct and disentangled target paths.
(a) Scaled positions over time without OT
(b) Scaled positions over time with OT
Figure 7: Ground Truth Token Velocities (scaled between -1 and 1). The ground truth velocities without OT (a) have a lot of crossing lines demonstrating instability in matching whereas the velocities with OT coupling (b) show almost straighter lines throughout the denoising process.
A.5 DDOT Training Algorithm
Algorithm 1 provides a concise overview of the DDOT training procedure. It details the steps involved in each training iteration:
  1. 1.
    Sampling a data instance $(x_{0},z_{0})$ from the dataset $\mathcal{D}$.
  2. 2.
    Sampling initial noisy positions $z_{T}\sim\mathcal{U}(-1,1)^{L}$ and a random timestep $t\sim\mathcal{U}(0,1)$.
  3. 3.
    Permuting the sampled $z_{T}$ to minimize the Optimal Transport cost with respect to the target positions $z_{0}$. This step is crucial for establishing the non-crossing paths within prompt and response segments.
  4. 4.
    Linearly interpolating between $z_{0}$ and the permuted $z_{T}$ to obtain the positions $z_{t}$ for the sampled timestep $t$.
  5. 5.
    Calculating the total loss $\mathcal{L}$ as a weighted sum of the token prediction loss $\mathcal{L}_{\text{tok}}$ and the position prediction loss $\mathcal{L}_{\text{pos}}$.
  6. 6.
    Updating the model parameters $\theta$ (which includes parameters for both $s_{\theta}$ and $v_{\theta}$) using the gradient of the total loss $\nabla\mathcal{L}$.
This process is repeated until convergence.
Algorithm 1 DDOT Training
1:Dataset $\mathcal{D}$, model $s_{\theta},v_{\theta}$, loss weight $\lambda$
2:while Training do
3:     $(x_{0},z_{0})\sim\mathcal{D}$; $z_{T}\sim\mathcal{U}(-1,1)^{L}$; $t\sim\mathcal{U}(0,1)$
4:     Permute $z_{T}$ to minimize OT cost
5:     Interpolate positions $z_{t}=(1-t)z_{0}+tz_{T}$
6:     $\mathcal{L}\leftarrow\mathcal{L}_{\text{tok}}+\lambda\mathcal{L}_{\text{pos}}$
7:     $\theta\leftarrow\text{Update}(\theta,\nabla\mathcal{L})$
8:end while

### v2

@@
 Abstract
-Discrete diffusion models are a new class of text generators that offer advantages such as bidirectional context use, parallelizable generation, and flexible prompting compared to autoregressive models.
-However, a critical limitation of discrete diffusion models is their inability to perform flexible-length or flexible-position text infilling without access to ground-truth positional data.
-We introduce DDOT (Discrete Diffusion with Optimal Transport Position Coupling), the first discrete diffusion model to overcome this challenge.
-DDOT jointly denoises token values and token positions, employing a novel sample-level Optimal Transport (OT) coupling.
-This coupling preserves relative token ordering while dynamically adjusting the positions and length of infilled segments, a capability previously missing in text diffusion.
-Our method is orthogonal to existing discrete text diffusion methods and is compatible with various pretrained text denoisers.
-Extensive experiments on text infilling benchmarks such as One-Billion-Word and Yelp demonstrate that DDOT outperforms naive diffusion baselines.
-Furthermore, DDOT achieves performance on par with state-of-the-art non-autoregressive models and enables significant improvements in training efficiency and flexibility.
+Discrete diffusion models are a new class of text generation models that offer advantages such as bidirectional context, parallelizable generation, and flexible prompting compared to autoregressive models.
+However, a critical limitation has been the inability to perform flexible-length or flexible-position text infilling without access to ground-truth positional data.
+We introduce DDOT [^1] Project page: https://andrew-zhang.github.io/ddot-page (Discrete Diffusion with Optimal Transport Position Coupling), a discrete diffusion model that overcomes this limitation by jointly denoising token values and token positions using a novel sample-level optimal transport coupling.
+This coupling preserves relative token ordering while dynamically adjusting the positions and lengths of infilled segments.
+DDOT is orthogonal to existing discrete text diffusion methods and is compatible with various pretrained text denoisers.
+On text-infilling benchmarks such as One-Billion-Word and Yelp, DDOT outperforms naive diffusion baselines and achieves performance on par with state-of-the-art non-autoregressive models, while improving training efficiency and prompting flexibility.
 Flexible-length Text Infilling for Discrete Diffusion Models
-Andrew Zhang, Anushka Sivakumar, Chiawei Tang, Chris Thomas Virginia Tech, Virginia, USA azhang42vt.edu
+Andrew Zhang    Anushka Sivakumar    Chiawei Tang    Chris Thomas Department of Computer Science Virginia Tech {azhang42,anushkas01,cwtang,christhomas}@vt.edu
 1 Introduction
-While autoregressive models are effective, they also involve highly sequential sampling, cannot take bidirectional context, and lead to constrained architectures by requiring a decoder mask.
+While autoregressive (AR) models are effective, they also involve highly sequential sampling, cannot use bidirectional context, and constrain architectures by requiring a decoder mask.
 In contrast, discrete diffusion models Lou et al.
-; Sahoo et al.
-(2024) can parallelize generation by denoising multiple tokens simultaneously, use bidirectional information, do not need a decoder mask, and, notably, allow for more controllable generation.
+(2024); Sahoo et al.
+(2024) can parallelize generation by denoising multiple tokens simultaneously, use bidirectional context, do not need a decoder mask, and allow for more controllable generation.
 Previous work Gat et al.
-; Shi et al.
-(2024) has demonstrated that discrete diffusion models are capable of arbitrary prompting locations, whereas autoregressive models are only capable of left-to-right text completion.
+(2024); Shi et al.
+(2024) has demonstrated that discrete diffusion models can handle prompts at arbitrary locations, whereas autoregressive models are only capable of left-to-right text completion.
 However, this advantage has a significant limitation: existing diffusion models cannot alter the distances between these prompt tokens.
 Consequently, existing text diffusion models cannot generate the ground-truth sample without access to the oracle positions of the prompt and infilled text.
-Figure 1: Our novel diffusion across token positions enables dynamic token movement for infilling.
-Unlike prior methods, DDOT learns to move mask tokens to appropriate locations, such as to the right of "brown," even if initially unmasked there.
-The OT coupling (colored lines) significantly simplifies this learning by drastically reducing possible permutations.
+Figure 1: Our diffusion across token positions enables dynamic token movement for infilling.
+Unlike prior methods, DDOT learns to move masked tokens to appropriate locations, such as to the right of “brown,” even if that position was not initially masked.
+The OT coupling (colored lines) simplifies this learning by drastically reducing the set of possible permutations.
 We solve this issue by enabling discrete diffusion models to learn where to move tokens.
-Specifically, we design a diffusion process that operates across token positions, allowing the model to vary the position and length of infilled spans.
+Specifically, we design a diffusion process that operates across token positions (in addition to token values), allowing the model to vary the positions and lengths of infilled spans (Figure 1).
 Furthermore, given the importance of token positioning in preserving semantic meaning He et al.
-, we incorporate sample-level OT (optimal transport) coupling to maintain relative token ordering throughout the diffusion process.
-Even minor positional changes can dramatically alter meaning, as seen in phrases like "The child’s green coat" and "The green child’s coat".
-DDOT’s OT coupling preserves this relative ordering throughout the diffusion process while supporting flexible-length text infilling.
-Our OT coupling prevents such swaps and drastically improves DDOT’s downstream performance across all studied benchmarks and metrics.
-Extensive experiments show that DDOT outperforms naive diffusion baselines and achieves on par performance with state-of-the-art non-autoregressive (NAR) models.
+(2020), we incorporate sample-level OT coupling Tong et al.
+(2024) to maintain relative token ordering throughout the diffusion process.
+Even minor positional changes can dramatically alter meaning, as seen in phrases like “The child’s green coat” and “The green child’s coat.” DDOT’s OT coupling preserves this relative ordering throughout the diffusion process while supporting flexible-length text infilling.
+Our OT coupling prevents such swaps and drastically improves DDOT’s training efficiency and downstream performance across all studied benchmarks and metrics.
+Extensive experiments show that DDOT outperforms naive diffusion baselines and achieves performance on par with state-of-the-art non-autoregressive (NAR) models.
 In summary, our contributions are as follows:
   • We propose DDOT, the first discrete text diffusion method for infilling arbitrary text sequences without ground-truth span lengths.
-  • We provide extensive experiments on DDOT that show it achieves outperforms diffusion baselines and achieves performance on par with state-of-the-art NAR models.
-  • We provide detailed ablations and visualizations that verify DDOT’s effectiveness for adjusting the position and length of infilled text spans and provide insights into our novel sample-level OT coupling.
-The OT coupling significantly outperforms naive diffusion across all tested benchmarks and metrics.
-Figure 2: DDOT learns to vary infilled span lengths and positions unlike prior fixed-position diffusion methods.
-(Left) We compute two separate intra-set OT couplings within the prompt positions and the response positions.
-This constraint drastically simplifies possible permutations.
-Right Given a timestep $t$, we predict the token and position.
+  • We provide extensive experiments on DDOT that show it outperforms diffusion baselines and achieves performance on par with state-of-the-art NAR models.
+  • We provide detailed ablations and visualizations that verify DDOT’s effectiveness in adjusting the positions and lengths of infilled text spans and provide insights into our novel sample-level OT coupling.
+The OT coupling significantly outperforms naive diffusion baselines across all tested benchmarks and metrics.
+Figure 2: DDOT learns to vary infilled span lengths and positions, unlike prior fixed-position diffusion methods.
+(Left) We compute two separate intra-set OT couplings within the prompt positions and the response positions, which drastically simplifies the set of possible permutations.
+(Right) Given a time step $t$, we predict the token and position.
 2 Related Work
-2.1 Lexicographically Constrained Generation
+2.1 Lexically Constrained Generation
 Constrained text generation has been explored through a variety of approaches, including AR and NAR methods Zhang et al.
@@
 (2020) enables flexible token generation through iterative insertion, though it still depends on sequential token prediction and multiple forward passes.
-AutoTemplate Iso (2024) approaches constrained text generation by simply feeding the prompt tokens into an encoder-decoder style model.
-CBART He (2021) extends the POINTER architecture by moving to an encoder-decoder model.
-Autoregressive methods, while effective for their specific use cases, inherit fundamental limitations: they require sequential generation that scales linearly with sequence length, and their causal attention masks prevent full utilization of bidirectional context during generation.
-Most critically, for text infilling tasks, these approaches struggle to simultaneously consider both past and future context when generating intermediate content Cao et al.
+AutoTemplate Iso (2024) approaches constrained text generation by feeding the prompt tokens into an encoder–decoder-style model.
+CBART He (2021) extends the POINTER architecture by moving to an encoder–decoder model.
+Autoregressive methods, while effective for their specific use cases, inherit fundamental limitations: they require sequential generation that scales linearly with sequence length, and their causal attention masks prevent the full use of bidirectional context during generation.
+Most critically, for text infilling tasks, these approaches struggle to consider both past and future contexts simultaneously when generating intermediate content Cao et al.
 (2023).
+Furthermore, to allow for flexible positions, autoregressive methods often regenerate the entire sequence instead of only inserting tokens, leading to wasted compute Iso (2024).
 2.2 Discrete Diffusion Models
 Discrete diffusion models offer an innovative approach to text generation, addressing key limitations of autoregressive methods Lou et al.
-; Ren et al.
+(2024); Ren et al.
 (2024); Gong et al.
@@
 (2024).
-These models denoise corrupted text sequences, enabling parallel updates of multiple tokens rather than the token-by-token process of autoregressive methods, reducing the number of forward passes.
-Additionally, their bidirectional nature allows them to leverage both past and future context in contrast to causal masking that constrains autoregressive models.
+These models denoise corrupted text sequences, enabling parallel updates of multiple tokens rather than the token-by-token process of autoregressive methods, thereby reducing the number of forward passes.
+Additionally, their bidirectional nature allows them to leverage both past and future context, in contrast to the causal masking that constrains autoregressive models.
 Early frameworks like D3PM Austin et al.
-adapted continuous diffusion to discrete tokens using Markovian corruption, while subsequent advances such as multinomial diffusion Hoogeboom et al.
-(2021) improved noise schedules and training efficiency.
+(2021) adapted continuous diffusion to discrete tokens by assigning probability that a token corrupts by either turning into a mask token or random token.
 Recent work on score-based discrete diffusion has further advanced the field by providing analytical solutions for the denoising process.
-Instead of directly modeling transition probabilities SEDD Lou et al.
-uses a score-based approach that learn the gradient of the log probability, which proves particularly effective for handling high-dimensional categorical data like text.
+Instead of directly modeling transition probabilities, SEDD Lou et al.
+(2024) uses a score-based approach that learns the gradient of the log-probability and introduces an entropy-based loss function, narrowing the gap to autoregressive models.
 However, despite these advantages, current discrete diffusion models face a significant limitation: they require fixed token positions throughout the generation process.
@@
 OT Villani et al.
-(2009) coupling has been well studied in image generation through theoretical foundations such as Rectified Flow, Liu et al.
-, which showed how continuous normalizing flows can be understood as OT paths.
-This framework demonstrated that probability mass can be efficiently transported through straight trajectories while preserving structural relationships, providing insights into how sequential data can be manipulated while maintaining order.
-Building on these insights, Tong et al.
-showed how OT can be incorporated into diffusion models to accelerate generation.
-By defining appropriate transport costs between positions, OT coupling ensures that dimensions within a sample move in straight, predictable paths.
-Although previous works incorporated OT into diffusion models with a focus on generation speed, our work leverages these properties to enable flexible-length text infilling.
-By coupling the discrete token values with continuous position variables through OT, we allow the model to optimize both token content and positioning simultaneously.
-This approach maintains the parallel generation and bidirectional context advantages of discrete diffusion while adding the ability to dynamically adjust sequence lengths.
-The OT coupling effectively serves as a bridge between the discrete token space and continuous position space, enabling natural language generation that respects both local token relationships and global sequence structure.
+(2009) coupling has been well studied in continuous diffusion.
+Tong et al.
+(2024) introduce minibatch couplings—either independent (random) or determined by OT—that regularize training and encourage near-straight trajectories.
+We adopt the term "coupling" from Tong et al.
+(2024) to denote a matching between an source and target distribution.
+By incentivizing straighter paths, minibatch OT coupling yields practical benefits such as faster sampling, a pattern also observed in OT-inspired methods like Rectified Flow (Liu et al., 2023) and Stochastic Interpolants (Albergo and Vanden-Eijnden, 2023).
+Our OT coupling differs in scope and granularity: we construct a sample-level coupling within each sequence, aligning token positions between an initial state and a target ordering.
+We instantiate this coupling via OT (and compare to independent coupling in ablations), which preserves relative order and guides tokens along predictable paths while still allowing the inter-set crossings needed for flexible-length infilling.
+Unlike minibatch couplings aimed primarily at accelerating continuous diffusion models, our coupling enables joint optimization of content and placement of tokens for discrete text diffusion, retaining parallel and bidirectional conditioning and permitting dynamic adjustment of span lengths and token spacing.
 3 Background: Masked Text Diffusion
 Discrete diffusion adapts the continuous nature of diffusion processes to text.
-The simplest and most performant version of discrete diffusion is masked diffusion.
-Rather than adding gaussian noise to continuous values such as pixels, masked text diffusion assigns a certain probability of masking tokens throughout the forward diffusion process.
-For the sake of this paper, masked diffusion can be seen as a masked language model (like BERT Devlin et al.
-(2019)) that works at gradually decreasing masking ratios to generate text.
+The simplest and most performant version of discrete diffusion is masked diffusion Lou et al.
+(2024); Shi et al.
+(2024); Sahoo et al.
+(2024).
+Rather than adding Gaussian noise to continuous values such as pixels, masked text diffusion assigns a probability of masking tokens throughout the forward diffusion process.
+For the purposes of this paper, masked diffusion can be seen as a masked language model (like BERT Devlin et al.
+(2019)) that operates at progressively decreasing masking ratios to generate text.
 Specifically, our text diffusion process follows Score Entropy Discrete Diffusion (SEDD) Lou et al.
-, modeling a score function across a support of $N$ states or token values.
-The forward diffusion process is given by a continuous time Markov chain (CTMC):
-$\frac{dp_{t}}{dt}=Q_{t}p_{t}\quad p_{0}\approx p_{\mathrm{data}}$ (1)
-$Q_{t}\in\mathbb{R}^{n\times n}$ denotes a transition matrix where the columns give the probability of transitioning from one discrete state (or token) to another.
-In masking diffusion models, the diagonal is $-1$, the last row is $1$, and all other elements are 00, moving probability mass from any token value to the last token in the vocabulary, the mask token.
-To extend the process to sequences rather than individual tokens, SEDD applies discrete diffusion independently to all tokens in a sequence.
-SEDD reverses the textual diffusion equation by modeling the score $s_{\theta}(x)_{y}$ of transitioning from token $x$ to token $y$ with a score entropy loss:
-${\scriptstyle\mathcal{L}_{\text{tok}}=\mathbb{E}_{(x,x_{0})}\Big{[}\sum_{y\neq x}w_{xy}\Big{(}s_{\theta}(x)_{y}-\frac{p(y|x_{0})}{p(x|x_{0})}\log s_{\theta}(x)_{y}\Big{)}\Big{]}}$ (2)
-where $w_{xy}$ weighs the importance of different states. Finally, to simulate the reverse diffusion process, we either take Euler steps or use an analytical denoiser Lou et al. .
+(2024), which models a score function over a support of $N$ states (token values).
+The forward diffusion process is described by a continuous-time Markov chain (CTMC):
+$\frac{dp_{t}}{dt}=Q_{t}p_{t}\quad p_{0}\approx p_{\mathrm{data}},$ (1)
+where $p_{t}\in\mathbb{R}^{N}$ is the column vector of state probabilities at time $t\in[0,T]$, $Q_{t}\in\mathbb{R}^{N\times N}$ is the (time-dependent) transition matrix (its columns sum to 0), and $p_{\mathrm{data}}\in\mathbb{R}^{N}$ is the empirical token distribution at $t=0$.
+In masked diffusion we include the mask token among the $N$ states and make it absorbing: each non-mask column has 1-1 on its diagonal and +1+1 in the mask row, while the mask column is all zeros, so columns sum to 0 and probability flows from non-mask tokens into the mask token.
+SEDD reverses the text diffusion dynamics by learning the score $s_{\theta}(x)_{y}$ of a transition from token $x$ to token $y$ via an entropy-based loss:
+$\mathcal{L}_{\text{tok}}=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0})}\sum_{y\neq x_{t}}Q_{t}(x_{t},y)\,\left(s_{\theta}(x_{t},t)_{y}-\right.\\[-2.0pt] \left.\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})}\log s_{\theta}(x_{t},t)_{y}+K\!\left(\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})}\right)\right)\,dt$ (2)
+where $x_{0}$ is the start token, $x_{t}$ is the token at time $t$ under the forward CTMC with marginal $p_{t|0}(\cdot|x_{0})$, $Q_{t}(x_{t},y)$ denotes the (row $x_{t}$, column $y$) entry of $Q_{t}$, and $K(a)=a(\log a-1)$ is a scalar regularizer; Finally, to simulate the reverse diffusion process, we either take Euler steps or use an analytical denoiser Lou et al.
+(2024).
 4 Approach
-Previous discrete diffusion models must know the ground truth positions of mask tokens to infill a sequence, as the positions of tokens are fixed throughout the entire diffusion process.
-To address this limitation, we introduce DDOT, a diffusion framework that jointly denoises discrete token values and continuous token positions, allowing for flexible infilling spans.
+Prior discrete diffusion models fix token positions during denoising and therefore require oracle positions to infill masked spans.
+We introduce DDOT, which jointly denoises discrete token values and continuous token positions so infilled spans can move and change length during generation.
+This enables flexible-length, flexible-position infilling from arbitrary prompts while preserving the parallel, bidirectional nature of discrete diffusion.
+Figure 2 summarizes DDOT.
+4.1 Token Value Diffusion
+We partition token values into three disjoint subsets $x=(x^{p},x^{r},x^{\mathrm{pad}})$, which are described as follows: (i) Prompt ($x^{p}$): these slots are clamped to their ground-truth tokens for all $t$; we never mask or resample them, and they serve purely as conditioning during denoising.
+(ii) Response ($x^{r}$): these are the real tokens to be infilled.
+They start at ground truth and follow the masked-diffusion forward process from section 3, progressively transitioning toward [mask] as $t$ increases; the reverse chain denoises them back to tokens.
+(iii) Pad ($x^{\mathrm{pad}}$): these slots hold the vocabulary token [pad] (distinct from [mask]), are not part of $x^{r}$, and exist only to maintain a fixed length $L$.
+Under the forward process (section 3), [pad] may corrupt to [mask]; in the reverse process, [mask] on these slots denoises back to [pad].
 4.2 Continuous Position Diffusion
-DDOT gradually transforms token positions from a simple initial distribution to their ground truth permutation.
-To achieve this, we denoise $l$ token positions $z_{t}\in[-1,1]^{l}$ in the continuous domain.
-The initial (limiting) distribution of positions is sampled from a uniform distribution, $z_{T}\sim\mathcal{U}(-1,1)^{L}$, where $L$ represents the maximum sequence length.
-The ground truth positions are defined as:
-$z_{0}=linspace\left(-\frac{l}{L},\frac{l}{L},l\right),$ (3)
-where $z_{0}$ is evenly spaced and scaled to match the length of the true sequence.
-This setup ensures that the position diffusion process captures the gradual transition from a simple prior to a structured output aligned with the token ordering and length.
-Similar to fixed position models, we scale by $l$ to provide information on the abosolute distances between tokens.
-Our preliminary studies reveal that naively performing diffusion on the token positions leads to poor performance.
-We observe that adding noise to token positions destroys the relative ordering of text and adds a combinatorially exploding number of permutations, problems not present in absolute-valued continuous domains such as image generation Ho et al.
-(2020); Karras et al.
-(2022).
-Since the relative ordering of text can drastically change its meaning and consequently abruptly destroy the signal across timesteps, it is crucial to fix token ordering across time.
-Unlike traditional diffusion processes that aim to completely destroy the data signal at the final timestep $T$, we provide the ordering of the prompt tokens at $T$ to the model.
-To maintain the prompt token ordering, we introduce a sample level OT coupling between the initial $z_{0}$ and final $z_{T}$ token positions of the diffusion process.
-Additionally, we do not add noise to the positions and instead use linear interpolation between $z_{0}$ and $z_{T}$.
-Following Liu et al. ; Albergo and Vanden-Eijnden , we model the forward position diffusion process with an ordinary differential equation (ODE)
-$dZ_{t}=v(Z_{t},t)dt$ (4)
-where $v(z_{t},t)$ is the velocity (rate of change) of the positions at time $t$.
-We partition $x$ and $z$ into prompt ($x^{p}$, $z^{p}$) and infilled response ($x^{r}$, $z^{r}$) subsets.
-The limiting positions $z_{T}$ are determined as the solution to the linear sum assignment problem, minimizing OT cost after scaling $z_{0}$in $[-1,1]$:
-$z^{*}_{T}=\pi(\frac{z_{0}L}{l},z_{T}),\quad\pi=\arg\min_{\pi^{\prime}\in S_{n}}\sum_{i=1}^{n}C_{i,\pi^{\prime}(i)}$ (5)
-where $C_{i,j}$ is the Euclidean distance between positions and $S_{n}$ is all possible permutations of length $n$.
-For prompt tokens, the optimal transport is balanced, so $|z^{p}_{T}|=|z^{p}_{0}|=l^{p}$ However, the number of infilled tokens in the limiting distribution is almost always larger than that of the ground truth $|z^{r}_{T}|=L-l^{p}\geq l-l^{p}=|z^{r}_{0}|$ To reconcile the different sizes of $z_{0}$ and $z_{T}$, we treat the remaining unassigned $L-l$ positions as pad tokens.
-Specifically, we set the ground truth tokens to the pad token and the ground truth positions to stationary paths before scaling: $z^{pad}_{0}=z^{pad}_{T}*\frac{l}{L}$.
-Importantly, these stationary paths maintain that the coupling between $z_{0}$ and $z_{T}$ achieves the optimal OT cost.
-This approach establishes a coupling within prompt and infilled tokens, ensuring that tokens in each subset maintain non-crossing paths, a key property of OT Tong et al.
-; Villani et al.
-(2009).
-While intra-set paths remain ordered to minimize transport cost, DDOT allows inter-set crossings (i.e., between prompt and infilled tokens), enabling flexible token positioning while preserving the relative order of prompt tokens.
-Path visualizations are available in A.4.
-Beyond prompt tokens, DDOT ensures that relative token order within all sets, prompt and infilled response, aligns with their ground truth at any timestep.
-This guarantees smooth transitions and avoids disruptions caused by abrupt positional swaps.
-Using OT, DDOT simplifies position manipulation, reducing unnecessary degrees of freedom.
-OT is traditionally computationally expensive, requiring iterative calculations on the CPU.
-However, our OT coupling is computationally efficient, addressing the large-scale data demands of generative text modeling.
-Previous methods approximate dataset-level OT through per-batch computations Tong et al.
-, typically relying on algorithms like the Sinkhorn method Cuturi (2013) and operating in high-dimensional spaces (e.g., image latents).
-In contrast, our OT coupling operates at the scalar level and limited to $L$ elements, corresponding to the model’s context size—a scale significantly smaller than the number of samples in a dataset.
-For prompt tokens, the OT coupling is balanced and can therefore be efficiently computed by simply sorting $z_{0}$ and $z_{T}$.
-This efficiency enables us to calculate exact sample-level OT, unlike the approximations required in previous dataset-level approaches.
-For training, we randomly sample a timestep $t$ and retrieve the corresponding positions in a simulation-free manner with linear interpolation $z_{t}=(1-t)z_{0}+tz_{T}$.
-The training objective is a weighted mean-squared error loss:
-${\mathcal{L}_{pos}(\theta)=\mathbb{E}_{(z,t)}\left[Q_{t}(x_{t},y)\|v_{\theta}(z_{t},t)-(z_{0}-z_{T})\|^{2}\right]}$ (6)
-We investigate two methods of initializing $z_{T}$ to begin generation.
-The first method is to sample $z_{T}\sim\mathcal{U}(-1,1)^{L}$, which we refer to as DDOT-random (DDOT-R).
-We then randomly select $l^{p}$ tokens from $z_{T}$ to serve as the positions for the prompt tokens.
-However, some areas of high density arise when randomly sampling which tend to map to pad tokens because there are less tokens in the corresponding ground truth region.
-Therefore, we propose DDOT-uniform (DDOT-U), which uniformly spaces out the prompt and infilled positions: $z^{p}_{T}=\text{linspace}(-1,1,l^{p})$ and $z^{r}_{T}=\text{linspace}(-1,1,l^{r})$.
-During the sampling process, the token paths tend to follow straight trajectories because the ground truth velocity, $z_{0}-z_{T}$, remains constant over time.
-Therefore, we find it sufficient to use Euler steps during sampling to predict the positions.
-Combined with $\tau$-leaping Lou et al.
-, the straight position paths allow for favorable compute-accuracy tradeoffs and result in fast sampling.
-The total loss $\mathcal{L}$ is a simple linear combination $\mathcal{L}=\mathcal{L}_{tok}+\lambda\mathcal{L}_{pos}$.
+To enable diffusion models to move masked tokens to regions that require infilling, DDOT also diffuses positions.
+Similar to token values, we partition positions into three disjoint subvectors $z_{t}=(z_{t}^{p},\,z_{t}^{r},\,z_{t}^{\mathrm{pad}})$ for prompt, response, and pad positions.
+Let $\ell=|x^{p}|+|x^{r}|$ be the number of real tokens, with per-subset counts $\ell^{p}=|x^{p}|$ and $\ell^{r}=|x^{r}|=\ell-\ell^{p}$; there are $L$-$\ell$ pad slots.
+The noisy positions are sampled i.i.d.
+as $z_{T}\sim\mathcal{U}(-1,1)^{L}$ (so $z_{T}^{p},z_{T}^{r},z_{T}^{\mathrm{pad}}$ are slices of a single length-$L$ vector).
+We define the ground-truth positions for the real (non-pad) tokens as the length-$\ell$ subvector over the union of prompt and response indices, $z_{0}^{p\cup r}\in\mathbb{R}^{\ell}$ (equivalently $z_{0}^{p\cup r}=(z_{0}^{p},z_{0}^{r})$), evenly spaced and scaled to the true length:
+$z^{p\cup r}_{0,i}=\frac{\ell}{L}\,\frac{2i-(\ell-1)}{\max\{1,\ell-1\}},\qquad i=0,\dots,\ell-1.$ (3)
+so real tokens occupy the sub-interval $\ell/L,\,\ell/L]\subset[-1,1]$.
+Linear position paths.
+Following Albergo and Vanden-Eijnden (2023) and Liu et al. (2023), positions follow straight, noise-free interpolation:
+$z_{t}\;=\;(1-t)\,z_{0}+t\,z_{T}.$ (4)
+Pad tokens.
+Pad tokens are part of the ground-truth vocabulary (distinct from [mask]). Pad positions remain stationary up to the same global scaling:
+$z^{\mathrm{pad}}_{0}\;=\;\frac{\ell}{L}\,z^{\mathrm{pad}}_{T}$ (5)
+Together with the $z^{p\cup r}$ construction in Equation 3, this induces the full length-$L$ slot vector $z_{0}=(z_{0}^{p},z_{0}^{r},z_{0}^{\mathrm{pad}})$.
+4.3 Positional Optimal Transport Coupling
+Our preliminary experiments (Table 4) show that naively diffusing continuous token positions performs poorly.
+It induces a combinatorial explosion of permutations, weakens supervision from the ground truth, and leads to invalid infilling results.
+We therefore use sample-level optimal-transport (OT) coupling.
+In particular, we find an OT coupling independently within each set $z^{p}$ and $z^{r}$.
+Our sample-level OT coupling has the following benefits:
+(a) Reduces the permutation space.
+OT yields order-preserving, non-crossing trajectories within each set (prompt or response), drastically shrinking the search over permutations.
+Concretely, for either set $S\in\{p,r\}$ and indices $1\leq i<j\leq\ell^{S}$ in that set’s ground-truth order, we have $z_{t}^{S,(i)}\leq z_{t}^{S,(j)}$ for all $t\in[0,1]$, so within-set crossings cannot occur (while inter-set prompt–response crossings remain allowed), avoiding the combinatorial explosion from arbitrary permutations.
+(b) Preserves supervision signal.
+Small word-order changes can flip semantics; The OT coupling between $z_{T}$ and $z_{0}$ preserves relative order throughout diffusion and thus maintains a stable learning signal from the ground-truth sequence rather than forcing the model to infer order from noisy, permuted trajectories.
+(c) Prevents invalid permutations.
+In infilling tasks, the prompt’s relative order is part of the conditioning; any output that reorders prompt tokens is an invalid infilling result.
+We feed the prompt tokens at the terminal state $T$ in their given order.
+With OT coupling, the forward $0\!\to\!T$ paths are order-preserving within the prompt set (non-crossing), so the reverse diffusion returns the same within-set order at $t=0$.
+Without coupling, trajectories can cross and swap prompt order, yielding invalid infills.
+Assignment formulation (per set).
+We compute the optimal transport coupling within prompt and non-prompt sets.
+From the noisy positions $z_{T}\in\mathbb{R}^{L}$, sample $\ell^{p}$ entries to form the prompt slice $z_{T}^{p}$; the remaining $L-\ell^{p}$ entries form the non-prompt slice $z_{T}^{\mathrm{np}}$.
+We compare ground-truth positions (rescaled to [1,1][-1,1]) directly to these slices.
+Prompt OT (balanced). One-to-one match: each of the $\ell^{p}$ prompt positions at $t=0$ is assigned to exactly one sampled prompt position at $t=T$, minimizing total displacement:
+$\pi^{p}=\operatorname*{arg\,min}_{\pi\in S_{\ell^{p}}}\sum_{i=1}^{\ell^{p}}\Bigl|\tfrac{L}{\ell}\,z^{p}_{0,i}-z^{p}_{T,\pi(i)}\Bigr|,\quad\tilde{z}_{T}^{p}=z_{T}^{p}[\pi^{p}].$ (6)
+Where $S_{\ell^{p}}$ is the set of permutations of $\ell^{p}]$ and, for any vector $v$, $v[\pi]$ denotes reindexing by $\pi$: $(v[\pi])_{i}\coloneqq v_{\pi(i)}$.
+Response OT (unbalanced).
+Injective match: the $\ell^{r}$ response positions at $t=0$ are assigned into the non-prompt pool $z_{T}^{\mathrm{np}}$ of size $L-\ell^{p}$, each slot used at most once; unassigned slots become pads:
+$\pi^{r}=\operatorname*{arg\,min}_{\pi\in\Pi_{\ell^{r}}^{\,L-\ell^{p}}}\sum_{i=1}^{\ell^{r}}\Bigl|\tfrac{L}{\ell}\,z^{r}_{0,i}-z^{\mathrm{np}}_{T,\pi(i)}\Bigr|,\quad\tilde{z}_{T}^{r}=z_{T}^{\mathrm{np}}[\pi^{r}].$ (7)
+Here $S_{\ell^{p}}$ is the set of all permutations of $\ell^{p}]$, and $\Pi_{\ell^{r}}^{\,L-\ell^{p}}$ is the set of injective maps $\ell^{r}]\hookrightarrow[L-\ell^{p}]$.
+Pads (leftover positions). The $L$-$\ell$ non-prompt positions not selected by $\pi^{r}$ constitute $z_{T}^{\mathrm{pad}}$. Pads are excluded from OT and follow the stationary path in Equation 5.
+After solving the assignments, we reorder the sampled positions and use the reordered slices for interpolation:
+$z_{t}^{S}\;=\;(1-t)\,z_{0}^{S}+t\,\tilde{z}_{T}^{S},\qquad S\in\{p,r\}.$ (8)
+This per-set OT coupling induces order-preserving, non-crossing trajectories within prompt and response (by construction of $\tilde{z}_{T}^{S}$ and linear paths), while allowing inter-set crossings between prompt and response.
+See subsection A.4 for visualizations.
+Computation.
+All couplings are computed per sequence in one dimension, using absolute distance between rescaled ground-truth positions and sampled positions.
+This choice lets us avoid forming dense cost matrices that are typically used when calculating OT and enables simple, fast implementations that scale favorably in the sequence length.
+Prompt OT computation.
+For the prompt set we must match $\ell^{p}$ ground-truth positions $z_{0}^{p}$ to $\ell^{p}$ sampled positions $z_{T}^{p}$ in a one-to-one manner.
+In 1D with convex costs, the optimal assignment is equivalent to sorting both vectors and pairing by index.
+Therefore, we sort $z_{0}^{p}$ and $z_{T}^{p}$ once and obtain the permutation $\pi^{p}$ that maps each sorted index in $z_{0}^{p}$ to the corresponding rank in $z_{T}^{p}$.
+This procedure achieves the exact minimum from Equation 6, runs in $O(\ell^{p}\log\ell^{p})$ time, and does not materialize a cost matrix.
+Response OT computation.
+We sort $a=z_{0}^{r}$ and $b=z_{T}^{\mathrm{np}}$ in ascending order, then compute the injective, order-preserving match with a standard global-alignment dynamic programming solution (Needleman and Wunsch, 1970).
+This exactly minimizes Equation 7 and runs in $O(\ell^{r}(L-\ell^{p}))$ time and $O(L-\ell^{p})$ memory.
+Complexity and practical overhead.
+The total per-sequence cost is $O(\ell^{p}\log\ell^{p}+\ell^{r}(L-\ell^{p}))$ time and $O(L)$ memory.
+For the context sizes considered ($L\leq 1024$), the coupling time is negligible compared to a single denoiser forward pass.
+In practice we precompute several OT couplings in advance on the CPU and stream the dataset to the GPU, leading to negligible overhead, as seen in subsection 5.3
+4.4 Training Objective
+At each update we draw a time $t\sim\mathcal{U}(0,1)$, sample noisy positions $z_{T}\sim\mathcal{U}(-1,1)^{L}$, and form the matched targets $\tilde{z}_{T}=(\tilde{z}_{T}^{p},\tilde{z}_{T}^{r},z_{T}^{\mathrm{pad}})$ by solving Equation 6Equation 7 (pads follow Equation 5).
+We then construct the per-set paths (Equation 8).
+Independently, for token values we follow SEDD’s forward process (Sec.
+3) to obtain $x_{t}$ from $x_{0}$ at the same $t$.
+Position loss.
+We supervise the position head (Figure 2) on all $L$ slots with a simple MSE toward the straight-path target:
+$\mathcal{L}_{\mathrm{pos}}(\theta)\;=\;\mathbb{E}_{t}\!\left[\,\big\|\,v_{\theta}(z_{t},x_{t},t)\;-\;\bigl(z_{0}-\tilde{z}_{T}\bigr)\big\|_{2}^{2}\right].$ (9)
+Token loss.
+For tokens we use the SEDD score-entropy objective from section 3, denoted $\mathcal{L}_{\mathrm{tok}}(\theta)$.
+Total loss.
+The final objective is a weighted sum
+$\mathcal{L}(\theta)\;=\;\mathcal{L}_{\mathrm{tok}}(\theta)\;+\;\lambda\,\mathcal{L}_{\mathrm{pos}}(\theta),$ (10)
+with $\lambda$ controlling the relative weight of position supervision.
+Initialization: DDOT-R vs. DDOT-U.
+During inference we manipulate how we sample $z_{T}$.
+By default (DDOT-R), we randomly sample a full length-$L$ terminal vector $z_{T}\sim\mathcal{U}(-1,1)^{L}$, reserve any $\ell^{p}$ entries as the prompt slice $z_{T}^{p}$.
+However, random sampling can produce high-density regions that tend to map to pad tokens because there are fewer tokens in the corresponding ground-truth region.
+To mitigate this, DDOT-U places terminal positions on uniform grids for each set:
+$z^{p}_{T,i}\;=\;\frac{\,2i-(\ell^{p}-1)\,}{\max\{1,\;\ell^{p}-1\}},$ $i=0,\dots,\ell^{p}-1,$ (11)
+$z^{np}_{T,i}\;=\;\frac{\,2i-(\ell^{np}-1)\,}{\max\{1,\;\ell^{np}-1\}},$ $i=0,\dots,\ell^{np}-1.$
+Uniform spacing avoids random clustering, reduces pad spillover, and yields more stable placements.
+Reverse-time updates.
+During sampling, positions follow near-straight trajectories because the learned velocity field approximates the constant ground-truth direction $z_{0}-\tilde{z_{T}}$.
+We therefore update positions with a simple Euler step at each $\tau$-leap of the SEDD reverse process, while keeping prompts clamped and pads stationary (Equation 5).
 4.5 Simultaneous Text & Position Diffusion
 DDOT performs discrete text and continuous position diffusion simultaneously, as these processes operate independently in continuous time.
-We therefore predict both token value scores and position velocities in a single forward pass.
-This independence also enables simulation-free training by independently sampling token and position states at arbitrary timesteps.
-We summarize the training procedure in Appendix A.5
+We therefore predict both token-value scores and position velocities in a single forward pass.
+This independence also enables simulation-free training by sampling token and position states independently at arbitrary time steps.
+We summarize the training procedure in algorithm 1.
+Algorithm 1 DDOT Training
+1:batch $(x_{0},z_{0})$; loss weight $\lambda$
+2:Sample $t\sim\mathcal{U}(0,1)$
+3:Build terminal positions: sample $z_{T}\sim\mathcal{U}(-1,1)^{L}$ and split $(z_{T}^{p},z_{T}^{\mathrm{np}})$, or use uniform grids from Equation 11
+4:Compute within-set OT: $\pi^{p}$ via 1D sort-match (Equation 6); $\pi^{r}$ via injective matching (Equation 7); set $\tilde{z}_{T}^{p}=z_{T}^{p}[\pi^{p}]$, $\tilde{z}_{T}^{r}=z_{T}^{\mathrm{np}}[\pi^{r}]$, leftovers $\to z_{T}^{\mathrm{pad}}$
+5:Form paths $z_{t}^{S}=(1-t)z_{0}^{S}+t\,\tilde{z}_{T}^{S}$ for $S\in\{p,r\}$; set pads via Equation 5
+6:Sample token states $x_{t}$ from the SEDD forward at time $t$
+7:Predict $s_{\theta}(x_{t},t)$ and $v_{\theta}(z_{t},x_{t},t)$
+8:Compute $\mathcal{L}_{\mathrm{tok}}$ (Equation 2) and $\mathcal{L}_{\mathrm{pos}}$ (Equation 9)
+9:Update $\theta$ on $\mathcal{L}_{\mathrm{tok}}+\lambda\,\mathcal{L}_{\mathrm{pos}}$ (Equation 10)
 4.6 Implementation Details
-Due to compute constraints, we implement our framework on SEDD Lou et al.
-.
-Our flexible position diffusion component is orthogonal to existing text diffusion methods and minimally alters the permutation of text tokens for infilling.
-This design allows DDOT to integrate seamlessly with various pretrained text diffusion models.
 We extend SEDD, which is based on the Diffusion Transformer architecture Peebles and Xie (2023), with two additional modules.
-First, we introduce a learnable type embedding applied directly after the token embedding lookup.
-This embedding indicates whether a token is part of the prompt or the masked response ($x\in x^{p}$ or $x\in x^{r}$), which is critical for assigning each token to the correct OT flow.
+First, we introduce a learnable type embedding applied directly after the token-embedding lookup.
+This embedding indicates whether a token is part of the prompt or the masked response ($x_{i}\in x^{p}$ or $x_{i}\in x^{n}p$), which is critical for assigning each token to the correct OT flow.
 Second, we add a linear head at the end of the Diffusion Transformer to compute $v_{\theta}(z_{t},t)$.
-To incorporate continuous positional information, we scale $z_{t}$ from the range $[-1,1]$ to match the context length of the original pretrained model (1024).
+The model architecture is available in Figure 2.
+To incorporate continuous positional information, we scale $z_{t}$ from [1,1][-1,1] to match the context length of the original pretrained model (1024).
 We then use Rotary Position Embeddings Su et al.
 (2024), a standard technique in discrete diffusion models.
-Implementation details can be found in subsection A.3.
+Implementation details can be found in Subsection A.3.2.
 5 Experiments
 5.1 Experimental Setup
-Datasets    We evaluate our approach on the One-Billion-Words and Yelp, following the preprocessing steps outlined in prior works on infilling and lexicographically constrained generation Miao et al.
+Datasets   We evaluate our approach on the One-Billion-Word and Yelp datasets, following the preprocessing steps outlined in prior work on infilling and lexically constrained generation Miao et al.
 (2019); Zhang et al.
@@
 These datasets consist of examples with $1$ to $6$ keywords that must be infilled while maintaining their relative order to generate coherent sentences.
-In addition to randomly masking positions, we also introduce the "block" masking method that masks a single continuous chunk of text 00 to $L/2$ tokens long (32 for one-billion-word and yelp and 512 for CodeParrot).
+In addition to randomly masking positions, we also introduce a block masking method that masks a single continuous span of text ranging from 0 to $L/2$ tokens (32 for One-Billion-Word and Yelp; 512 for CodeParrot).
 Finally, we apply the aforementioned masking methods to the Python subset of the CodeParrot dataset.
-Table 1 illustrates examples of this lexicographically constrained generation task.
+Table 1 illustrates examples of this lexically constrained generation task.
 Keywords: earned, cents, share, costs
@@
 Table 1: Example generations for the keywords-to-sentence generation on One-Billion-Word and Yelp.
-Model One-Billion-Word-Block Yelp-Block CodeParrot-Block CodeParrot-Random
+Model One Billion Word–Block Yelp-Block CodeParrot-Block CodeParrot-Random
 B2/B4 N2/N4 M SR B2/B4 N2/N4 M SR B2/B4 CB SR B2/B4 CB SR
@@
 Table 2: DDOT outperforms diffusion baselines on standard sequences (0-32 prompt tokens). Metrics are BLEU (B2, B4), NIST (N2, N4), METEOR (M), and Success Rate (SR). Top scores are bolded.
-Model One-Billion-Word-Random Yelp-Random
+Model One Billion Word–Random Yelp-Random
 B2/B4 N2/N4 M SR B2/B4 N2/N4 M SR
@@
 Traditional Models
-   SeqBF Mou et al. (2016) 4.4/0.7 0.62/0.62 7.0 $\leq 100.$ 6.9/2.1 0.52/0.53 8.7 $<100.$
+ SeqBF Mou et al. (2016) 4.4/0.7 0.62/0.62 7.0 $\leq 100.$ 6.9/2.1 0.52/0.53 8.7 <100.<100.
  CGMH Miao et al. (2019) 9.9/3.5 1.15/1.17 13.1 100. 12.3/4.6 1.41/1.45 14.6 $\leq 100.$
@@
 Figure 3: Success rate on block datasets. LC and PoP increasingly generate invalid responses (missing or swapping prompt tokens) as the number of prompt tokens grows.
-Training Details    To align the position prediction modules, we first finetune SEDD with the added modules on FineWeb-Edu Penedo et al.
+Training Details   To align the position-prediction modules, we first fine-tune SEDD with the added modules on FineWeb-Edu Penedo et al.
 (2024).
-Afterwards, we further finetune on One-Billion-Word and Yelp datasests.
-For simplicity, we always keep all parameters unfrozen and simultaneously optimize both $\mathcal{L}_{tok}$ and $\mathcal{L}_{pos}$.
+Afterward, we further fine-tune on the One-Billion-Word and Yelp datasets.
+For simplicity, we keep all parameters unfrozen and optimize $\mathcal{L}_{\text{tok}}$ and $\mathcal{L}_{\text{pos}}$ simultaneously.
 In line with SEDD, we train our model in two configurations: small (90M non-embedding parameters) and medium (320M non-embedding parameters).
-DDOT-medium is on the same scale as CBART (406M parameters), and AutoTempalate-base (220M parameters).
-Following SEDD, we use the AdamW optimizer with a learning rate of 3e-5.
-We set $\lambda=10$ when using scaling.
-For each experiment, we either use 48 L40s (48 GB) GPUs, 80 A30 (24GB) GPUs, or 8 A100 (80GB) GPUs.
+DDOT-medium is on the same scale as CBART (406M parameters) and AutoTemplate-base (220M parameters).
+Following SEDD, we use the AdamW optimizer with a learning rate of $3\times 10^{-5}$.
+We set $\lambda=10$ for position-loss weighting.
+For each experiment, we use either $48\times$ L40 (48 GB), $80\times$ A30 (24 GB), or $8\times$ A100 (80 GB) GPUs.
 Baselines   We compare our method against strong autoregressive (AR) and non-autoregressive (NAR) baselines.
-AutoTemplate Iso (2024), the state-of-the-art AR model, leverages the T5 Raffel et al.
-(2020) family of pretrained models.
-Specifically, AutoTemplate parses the lexicographically constrained generation task into a template that is autoregressively generated from left to right.
-The previous state-of-the-art NAR method, CBART He (2021), is built upon the BART Lewis et al.
-(2020) pretrained framework and iteratively inserts tokens into a sequence.
-We also introduce two diffusion-based models that follow the same training procedure as DDOT.
-Left Context (LC) concatenates all the prompt tokens to the left of the sequence and generates the response to the right of a separator token.
-Position Prediction (PoP) uses a SEDD model with a linear head that first predicts the positions of every token.
-Then, this sequence is fed through a finetuned fixed-position SEDD.
-Distribution Annealing    Many lexiocagphically constrained generation baselines including AutoTemplate and CBART use distribution annealing methods such as top-p, top-k, greedy sampling, and beam search.
-To provide a parallel to greedy decoding which always takes the top token probability, we anneal the distributions of our tokens values during sampling to only include the most probable token.
-Specifically, given the predicted probability of a certain token being mask $\hat{p}(x^{mask})$, we assign $1-\hat{p}(x^{mask})$ to the token value with the highest probability excluding the mask token.
-The rest of the token probabilities are set to 00.
-Greedy decoding in prior models (such as autoregressive) is deterministic, collapsing the tree of all generation paths into a single path.
-However, our annealing process maintains generation diversity (A.2) because the model must still sample from the annealed distribution with the top token value and the mask token.
-Whenever possible, we evaluate against the greedy decoding baseline.
-Metrics  Following prior works Miao et al.
+AutoTemplate Iso (2024), a state-of-the-art AR approach, leverages the T5 Raffel et al.
+(2020) family of pretrained models and parses the lexically constrained generation task into a template that is generated autoregressively from left to right.
+The previous state-of-the-art NAR method, CBART He (2021), is built on the pretrained BART framework Lewis et al.
+(2020) and iteratively inserts tokens into a sequence.
+We also introduce two diffusion-based baselines that follow the same training procedure as DDOT.
+Left Context (LC) concatenates all prompt tokens to the left of the sequence and generates the response to the right of a separator token.
+Position Prediction (PoP) uses a SEDD model with a linear head that first predicts the positions of every token; this sequence is then fed through a fine-tuned fixed-position SEDD.
+Distribution Annealing   Many lexically constrained generation baselines, including AutoTemplate and CBART, use distribution-annealing methods such as top-$p$, top-$k$, greedy decoding, and beam search.
+To provide a parallel to greedy decoding—which always takes the top token probability—we anneal the token-value distribution during sampling to include only the most probable non-mask token.
+Specifically, given the predicted probability that a token is the mask, $\hat{p}(x^{\text{mask}})$, we assign $1-\hat{p}(x^{\text{mask}})$ to the token value with the highest probability excluding the mask token, and set the remaining token probabilities to 0.
+Greedy decoding in prior models (e.g., AR) is deterministic, collapsing the tree of all generation paths into a single path.
+However, our annealing process maintains generation diversity (A.2.2) because the model still samples from the annealed distribution over the top token value and the mask token.
+Whenever possible, we evaluate against the greedy-decoding baseline.
+Metrics  Following prior work Miao et al.
 (2019); Zhang et al.
-(2020); Iso (2024), we evaluate on BLEU-2/4 Papineni et al.
-(2002), NIST-2/4 Doddington (2002), and METEOR-v1.5 Denkowski and Lavie (2014), and success rate (SR).
+(2020); Iso (2024), we evaluate BLEU-2/4 Papineni et al.
+(2002), NIST-2/4 Doddington (2002), METEOR-v1.5 Denkowski and Lavie (2014), and success rate (SR), which is the percentage of responses that are valid infilling results.
+A result is determined as invalid when any prompt tokens in the generated sequence do not appear in the same relative order that they were provided.
 5.2 Main Results
-We present lexicographically constrained generation results in Table 2.
-Our approach uses greedy annealing and is compared against greedy decoding wherever applicable, including the CBART greedy decoding baseline.
+We present lexically constrained generation results in Table 2.
+Our approach uses greedy annealing and is compared against greedy decoding wherever applicable, including the CBART greedy-decoding baseline.
 Our method achieves competitive performance with previous NAR models, approaching AR performance.
-Notably, our model achieves state of the art performance on most metrics among the diffusion baselines.
-Our method does well on block infilling, which may be more useful in real-world applications.
-Furthermore, we notice that DDOT scales well to longer sequences.
-Specifically, it frequently generates valid responses that include all the prompt words in the same relative ordering, as shown in the success rates (SR).
-In contrast, diffusion baselines quickly generate invalid responses as the number of prompt tokens increase (Figure 3, Table 2).
-However, in benchmarks with 6 or less prompt tokens, diffusion baselines maintain high SR (Table 3).
-This may be because the fixed-position models have room to correct generation when the ratio of prompt to response tokens is low.
-Table 3 Compares results with previous works in the lexicographically constrained generation line of work.
-Since pretrained diffusion models lag behind AR models, an issue not unique to DDOT, we elect to focus on NAR models.
-DDOT performs on par with the previous SOTA models, and achieves higher SR than all diffusion baselines.
-Although DDOT underperforms LC and PoP in some metrics, we argue that one-billion-word-random and yelp-random over-index on the unrealistic task of generating text with only $1-6$ randomly spaced tokens, and that Figure 3 shows the broader trend when DDOT is scaled to more prompt tokens.
+Notably, our model achieves state-of-the-art performance on most metrics among the diffusion baselines.
+Our method performs well on block infilling, which may be more useful in real-world applications.
+Furthermore, we observe that DDOT scales well to longer sequences: it frequently generates valid responses that include all prompt words in the same relative order, as reflected in the SR.
+In contrast, diffusion baselines quickly generate invalid responses as the number of prompt tokens increases (Figure 3, Table 2).
+However, in benchmarks with six or fewer prompt tokens, diffusion baselines maintain high SR (Table 3).
+This may be because fixed-position models have room to correct generation when the ratio of prompt to response tokens is low.
+Table 3 compares results with previous work in lexically constrained generation.
+Since pretrained diffusion models currently lag behind AR models—an issue not unique to DDOT—we focus on NAR models.
+DDOT performs on par with previous state-of-the-art models and achieves higher SR than all diffusion baselines.
+Although DDOT underperforms LC and PoP on some metrics, we argue that the One-Billion-Word–Random and Yelp–Random settings over-index on the unrealistic task of generating text from only 1166 randomly spaced tokens.
+Furthermore, Figure 3 shows the broader trend when DDOT is scaled to a larger number of prompt tokens.
 5.3 Analysis
-In this section we investigate the effect of random versus uniform position initialization, the inclusion of OT coupling, and the impact of varying the number of sampling steps.
-Figure 4: (a) and (b) exhibit Ground Truth Token Velocities.
-The velocities without OT (a) have a lot of crossing lines demonstrating instability in matching whereas the velocities with OT coupling (b) show almost straighter lines throughout the denoising process.
-(c) Performance tends to increase with sampling steps.
-Position Initialization    In Table 2 and Table 3 we also explore the difference between DDOT-R and DDOT-U.
-In DDOT-R, pad tokens tend to cluster in areas of high density because the OT finds no match for them.
-However, the pad tokens in DDOT-U tend to be evenly spaced out.
+In this section, we investigate the effect of random versus uniform position initialization, the inclusion of OT coupling, and the impact of varying the number of sampling steps.
+Figure 4: (a) and (b) show ground-truth token positions over time.
+Without OT (a), many line crossings indicate unstable permutations, whereas with OT coupling (b), trajectories are nearly straight throughout the denoising process.
+(c) Performance tends to increase with more sampling steps.
+Position Initialization   In Table 2 and Table 3, we also explore the difference between DDOT-R and DDOT-U.
+In DDOT-R, pad tokens tend to cluster in high-density regions because OT finds no matches for them; in DDOT-U, pad tokens tend to be evenly spaced.
 We find that DDOT-U consistently outperforms DDOT-R.
@@
 Models trained with OT coupling consistently outperform those using independent (random) coupling.
-We theorize that the OT coupling provides more signal about the ordering of tokens throughout the diffusion process.
-Specifically, the DDOT guarantees that the relative ordering of the prompt and generated tokens at any timestep is the same as the original ordering.
-In contrast, independent coupling requires the model to infer the original ordering of tokens—a challenging task given the numerous plausible orderings that can result from interspersed prompt tokens.
+We hypothesize that OT coupling provides a stronger signal about token ordering throughout the diffusion process.
+Specifically, DDOT guarantees that the relative ordering of the prompt and generated tokens at any time step matches the original ordering.
+In contrast, independent coupling requires the model to infer the original ordering—a challenging task given the numerous plausible orderings that can result from interspersed prompt tokens.
 OT? B2 B4 N2 N4 M
@@
 Table 4: Our OT coupling drastically improves preformance across all metrics. Ablation on OT coupling with small model size.
-Position Over Time    We qualitatively compare the ground truth token paths during training Figure 4.
-With OT coupling, token trajectories exhibit significantly fewer crossings, maintaining relative order throughout the generation process.
-In contrast, the independent coupling frequently permutes tokens.
-Visualizations of token paths during inference are available in subsection A.4
-Number of Sampling Steps    One advantage of diffusion models over autoregressive models is their ability to exchange compute for accuracy by varying the number of inference steps.
-Figure 4 shows how the number of sampling steps influences lexicographically constrained generation performance.
-As the number of sampling steps increases, performance also increases.
-Wall Time Analysis    We evaluate the inference speed of DDOT against the diffusion baselines on the One-Billion-Word dataset.
-Table 5 presents the wall-clock time per batch alongside BLEU-2 and BLEU-4 scores for an increasing number of sampling steps.
+Position Over Time   We qualitatively compare ground-truth token paths during training in Figure 4.
+With OT coupling, token trajectories exhibit significantly fewer crossings, maintaining relative order throughout the generation process; in contrast, independent coupling frequently permutes tokens.
+Visualizations of token paths during inference are available in subsection A.4.
+Number of Sampling Steps   One advantage of diffusion models over autoregressive models is the ability to trade compute for accuracy by varying the number of inference steps.
+Figure 4 shows how the number of sampling steps influences lexically constrained generation performance: as the number of sampling steps increases, performance improves.
+Wall-Time Analysis   We evaluate the inference speed of DDOT against the diffusion baselines on the One-Billion-Word.
+Table 5 presents the wall-clock time per batch alongside BLEU-2 and BLEU-4 scores for increasing numbers of sampling steps.
 DDOT demonstrates significantly better efficiency.
-For any given number of sampling steps, DDOT is not only faster than LC and competitive with PoP in terms of raw speed, but also achieves substantially higher BLEU scores.
+For any given number of sampling steps, DDOT is not only faster than LC and competitive with PoP in raw speed, but also achieves substantially higher BLEU scores.
 Notably, LC must regenerate prompt tokens and therefore requires up to double the input sequence length.
@@
 Table 5: DDOT achieves superior BLEU scores with faster inference times. Inference speed (seconds per batch) and BLEU scores on One-Billion-Word for varying numbers of sampling steps.
-The added modules to enable position prediction are lightweight, consisting of a linear head and two type embeddings.
-On the other hand, the LC baseline requires double the context length of DDOT because it must regenerate prompt tokens.
-The OT calculation is highly efficient, taking 16 minutes and 11 seconds on an Intel Xeon 8462Y+ 64 core processor for the 10 billion token subset of FineWeb EDU.
-In practice, we stream the dataset, caching several OT couplings in advance without needing to preprocess the OT.
-With caching, it takes 4 minutes and 30 seconds to do 1000 training steps on an L40 gpu with a batch size of 256.
-Without caching, it takes 4 minutes and 27 seconds, a negligible difference.
+The added modules that enable position prediction are lightweight, consisting of a linear head and two type embeddings.
+By contrast, the LC baseline requires double the context length of DDOT because it must regenerate prompt tokens.
+The OT calculation is highly efficient, taking 16 minutes, 11 seconds on an Intel Xeon 8462Y+ 64-core processor for the 10-billion-token subset of FineWeb-Edu.
+In practice, we stream the dataset, caching several OT couplings in advance without needing to preprocess them.
+With caching, it takes 4 minutes, 30 seconds to run 1,000 training steps on an L40 GPU with a batch size of 256.
+Without caching, it takes 4 minutes, 27 seconds—a negligible difference.
 6 Conclusion
-In this work we introduce DDOT, the first discrete diffusion model capable of flexible-length text infilling by jointly denoising token values and positions.
-By incorporating optimal transport coupling, DDOT preserves token order while enabling dynamic position adjustment, addressing limitations in existing text diffusion models.
-Our experiments show that DDOT outperforms diffusion baselines and matches SOTA non-autoregressive models.
-7 Limitations
+In this work, we introduce DDOT, the first discrete diffusion model capable of flexible-length text infilling by jointly denoising token values and positions.
+By incorporating optimal-transport coupling, DDOT preserves token order while enabling dynamic position adjustment, addressing limitations in existing text diffusion models.
+Our experiments show that DDOT outperforms diffusion baselines and matches state-of-the-art non-autoregressive models.
+Limitations
 DDOT inherits many drawbacks from previous discrete text diffusion methods Lou et al.
-; Austin et al.
-; Sahoo et al.
+(2024); Austin et al.
+(2021); Sahoo et al.
 (2024).
@@
 Since diffusion for text is an emerging field, this performance gap is expected.
-Similarly, the finetuned versions of these models (such as DDOT vs AutoTemplate) also exibit the same performance gap.
-Second, unlike autoregressive models, existing diffusion models cannot use a kv-cache to store previous activations.
-DDOT also inherits risks common between large-scale text models.
-8 Ethics Statement
-Artifacts  The artifacts that we used had public use licenses. Furthermore, we plan to release our code artifacts for public use after acceptance.
-Dataset Considerations  We use publicly available datasets that underwent safety checks such as FineWeb-EDU.
-Documentaion of Artifacts  Our work generates English text, and our codebase is primarily in Python.
+Similarly, the fine-tuned versions of these models (e.g., DDOT vs.
+AutoTemplate) exhibit the same gap.
+Second, unlike autoregressive models, existing diffusion models cannot use a KV cache to store previous activations.
+DDOT also inherits risks common to large-scale text models.
+Ethical Considerations
+Artifacts The artifacts we used have public-use licenses. Furthermore, we plan to release our code artifacts for public use after acceptance.
+Dataset Considerations We use publicly available datasets that underwent safety checks, such as FineWeb-Edu.
+Documentation of Artifacts Our work generates English text, and our codebase is primarily in Python.
 Use of AI Assistants We used AI assistants to help write our code and revise our paper.
-Package Usage  we use nltk for BLEU, NIST, METEOR, and N gram diversity. We use https://pypi.org/project/codebleu/ for codebleu.
+Package Usage We use NLTK for BLEU, NIST, METEOR, and $n$-gram diversity. We use https://pypi.org/project/codebleu/ for CodeBLEU.
+Acknowledgements
+We thank Virginia Tech’s Advanced Research Computing (ARC) for computational resources and technical support, and the anonymous reviewers for their constructive feedback.
 Appendix A Appendix
-This appendix provides supplementary material to support the findings and methodology presented in the main paper.
-It includes detailed model size comparisons, generation diversity analysis, hyperparameter sensitivity studies, visualizations of token path dynamics, and the DDOT training algorithm.
-A.1 Size Comparison
-Table 6 presents a detailed breakdown of performance metrics (BLEU-2/4, NIST-2/4, and METEOR) for different sizes of our DDOT model (small and medium) and its variants (DDOT-R and DDOT-U).
-These results are shown for the One-Billion-Word-Random and Yelp-Random benchmarks.
-This comparison allows for an assessment of model scalability and the impact of the $z_{T}$ initialization strategy (random vs.
-uniform) across different model capacities.
-The medium-sized DDOT-U variant generally demonstrates the strongest performance, highlighting the benefits of uniform initialization and increased model parameters.
+A.1 Overview
+This appendix provides additional results, diversity analyses, hyperparameter sensitivity, path visualizations, and short implementation notes for reproducibility.
+We reuse notation from the main text and make one explicit addition used below:
+$\ell^{\mathrm{np}}\;=\;L-\ell^{p}$
+denoting the number of non-prompt slots (response ++ pads). This matches the usage of $z_{T}^{\mathrm{np}}$ in the main text.
+A.2 Additional Results
+A.2.1 Size Comparison
+Table 6 reports BLEU-2/4, NIST-2/4, and METEOR for ddot small/medium and for ddot-R vs.
+ddot-U on One-Billion-Word–Random and Yelp–Random.
+Uniform $z_{T}$ initialization (ddot-U) consistently outperforms random (ddot-R), and scaling to the medium model further improves most metrics.
 Model One-Billion-Word-Random Yelp-Random
@@
 Table 6: Results on One-Billion-Word and Yelp. Metrics are BLEU (B2, B4), NIST (N2, N4), and METEOR (M). Best results are highlighted in bold.
-A.2 Generation Diversity
-Table 7 provides an evaluation of generation diversity for DDOT and baseline models on the One-Billion-Word-Random and Yelp-Random datasets.
-Diversity is measured using D2 and D4 metrics, which represent the percentage of unique bi-grams and 4-grams in the generated outputs, respectively.
-These metrics are crucial for assessing the variety and richness of the generated text beyond simple n-gram overlap with references.
-The results show how DDOT variants compare to other baselines like CBART, LC, and PoP in this regard, with CBART exhibiting high diversity scores.
-DDOT models offer competitive diversity, particularly DDOT-U on the One-Billion-Word dataset.
+A.2.2 Generation Diversity
+Table 7 compares diversity using D2/D4 (unique bigrams/four-grams). ddot-U yields competitive diversity relative to strong NAR baselines while maintaining ordering constraints.
 Model One-Billion-Word-Random Yelp-Random
@@
 Table 7: Results on One-Billion-Word and Yelp dataset for generation diversity. Metrics include Diversity (D2, D4). Best results are highlighted in bold.
-A.3 Hyperparameter Sensitivity Analysis
-This section details the hyperparameters explored and used in our experiments, providing transparency and aiding reproducibility.
-A.3.1 $\lambda$ values
-The choice of the loss weight $\lambda$ is critical as it balances the contribution of the token prediction loss ($\mathcal{L}_{\text{tok}}$) and the position prediction loss ($\mathcal{L}_{\text{pos}}$).
-Table 8 investigates the impact of different $\lambda$ values on model performance metrics (B2, B4, N2, N4, M) for the One-Billion-Word dataset.
-The results demonstrate the model’s performance across a range of $\lambda$ values, indicating a degree of robustness.
-For the experiments reported in this table, a $\lambda$ value around 10 appears to offer a good balance.
+A.3 Hyperparameter Sensitivity and Training Details
+A.3.1 Loss Weight $\lambda$
+We sweep $\lambda$ to balance $\mathcal{L}_{\text{tok}}$ and $\mathcal{L}_{\text{pos}}$.
+Results in Table 8 show stable performance across a range; $\lambda\!\approx\!10$ is a good default used in the main results.
 $\lambda$ B2 B4 N2 N4 M
@@
 A.3.2 Training Hyperparameters
-Table 9 lists the primary hyperparameters used for training the DDOT models presented in the main paper.
-These settings, including batch size variations (dependent on GPU resources like A30, L40, or A100), the $\lambda$ value used with scaling (300), weight decay, learning rate, optimizer choice (AdamW), and its $\beta$ parameters, are provided for reproducibility.
-These parameters were selected based on preliminary experiments and common practices in training diffusion models.
-Most parameters were selected based on SEDD’s values.
-$\lambda$ was emperically determined from small test runs.
+[…]
 Hyperparameter Value
@@
 Weight decay 0.0
-Learning rate $43e-$ 4
+Learning rate 3e43e-4
 Optimizer AdamW
@@
 A.4 Path Visualizations
-The following figures provide qualitative insights into the token path dynamics during the diffusion process.
-These visualizations help to understand the behavior of token positions during inference (sampling) and the nature of the ground truth paths the model learns during training, both with and without Optimal Transport (OT) coupling.
-Figure 6 visualizes token trajectories during the inference (sampling) process without OT coupling, shown for both scaled (a) and unscaled (b) positions across a number of sampling steps.
-These plots illustrate how token paths can become less orderly and may curve or cross significantly towards the end of the generation process when OT is not employed to maintain relative order, potentially leading to permutation errors.
-(a) Scaled positions over time without OT
-(b) Unscaled positions over time without OT
-Figure 5: Ablations on number of sampling steps. Token paths tend to curve towards the end of generation without OT coupling.
-Figure 6, in contrast, shows token trajectories during inference (sampling) *with* OT coupling, again for scaled (a) and unscaled (b) positions.
-These paths are visibly straighter and more ordered throughout the generation process.
-This demonstrates the effectiveness of OT in maintaining relative token order and guiding tokens towards their target positions more directly, leading to more stable and accurate generation.
-(a) Scaled positions over time with OT
-(b) Unscaled positions over time with OT
-Figure 6: Ablations on number of sampling steps. Token paths tend to be straighter towards the end of generation with OT coupling.
-Figure 7 depicts the ground truth token paths between $z_{T}$ (initial noisy state) and $z_{0}$ (target state) during training, scaled between -1 and 1.
-Subfigure (a) shows these paths *without* OT coupling; here, $z_{T}$ represents a more arbitrary mapping from initial random positions to $z_{0}$, leading to many crossing paths.
-This illustrates the complex, potentially tangled mapping the model would need to learn.
-Subfigure (b) shows the paths *with* our sample-level OT coupling.
-In this case, $z_{T}$ is aligned with $z_{0}$ via optimal transport (within prompt and response sets), resulting in significantly straighter, non-crossing (within-set) trajectories.
-This comparison highlights how OT simplifies the learning objective for the positional diffusion component by providing more direct and disentangled target paths.
-(a) Scaled positions over time without OT
-(b) Scaled positions over time with OT
-Figure 7: Ground Truth Token Velocities (scaled between -1 and 1).
-The ground truth velocities without OT (a) have a lot of crossing lines demonstrating instability in matching whereas the velocities with OT coupling (b) show almost straighter lines throughout the denoising process.
-A.5 DDOT Training Algorithm
-Algorithm 1 provides a concise overview of the DDOT training procedure. It details the steps involved in each training iteration:
-  1. 1.
-    Sampling a data instance $(x_{0},z_{0})$ from the dataset $\mathcal{D}$.
-  2. 2.
-    Sampling initial noisy positions $z_{T}\sim\mathcal{U}(-1,1)^{L}$ and a random timestep $t\sim\mathcal{U}(0,1)$.
-  3. 3.
-    Permuting the sampled $z_{T}$ to minimize the Optimal Transport cost with respect to the target positions $z_{0}$.
-This step is crucial for establishing the non-crossing paths within prompt and response segments.
-  4. 4.
-    Linearly interpolating between $z_{0}$ and the permuted $z_{T}$ to obtain the positions $z_{t}$ for the sampled timestep $t$.
-  5. 5.
-    Calculating the total loss $\mathcal{L}$ as a weighted sum of the token prediction loss $\mathcal{L}_{\text{tok}}$ and the position prediction loss $\mathcal{L}_{\text{pos}}$.
-  6. 6.
-    Updating the model parameters $\theta$ (which includes parameters for both $s_{\theta}$ and $v_{\theta}$) using the gradient of the total loss $\nabla\mathcal{L}$.
-This process is repeated until convergence.
-Algorithm 1 DDOT Training
-1:Dataset $\mathcal{D}$, model $s_{\theta},v_{\theta}$, loss weight $\lambda$
-2:while Training do
-3:     $(x_{0},z_{0})\sim\mathcal{D}$; $z_{T}\sim\mathcal{U}(-1,1)^{L}$; $t\sim\mathcal{U}(0,1)$
-4:     Permute $z_{T}$ to minimize OT cost
-5:     Interpolate positions $z_{t}=(1-t)z_{0}+tz_{T}$
-6:     $\mathcal{L}\leftarrow\mathcal{L}_{\text{tok}}+\lambda\mathcal{L}_{\text{pos}}$
-7:     $\theta\leftarrow\text{Update}(\theta,\nabla\mathcal{L})$
-8:end while
+We visualize token-position trajectories during both inference (what the model actually does at test time) and training (the straight paths used as supervision).
+We show two coordinate systems: scaled to [1,1][-1,1] (used in losses; cf.
+Equation 3) and unscaled in the model’s native context length (for intuition about absolute movement).
+Unless noted, we split by set $S\!\in\!\{p,r\}$ and index tokens within a set by their ground-truth order.
+Plot Explanation.
+Let $\{t_{u}\}_{u=0}^{K}$ be the reverse-time grid (left-to-right on the $x$-axis), with $t_{0}=T$ and $t_{K}=0$.
+For each token index $i$ in set $S$, we plot its 1D position $z^{S,(i)}_{t_{u}}$ across $u$.
+Colors denote the token’s state (masked vs.
+unmasked) at each step.
+During inference, $z_{t_{u+1}}$ is obtained by an Euler update using the learned velocity $v_{\theta}(z_{t_{u}},x_{t_{u}},t_{u})$ while clamping prompts and keeping pads stationary (Equation 5).
+During training, targets follow straight paths $z_{t}^{S}=(1-t)z_{0}^{S}+t\,\tilde{z}_{T}^{S}$ from Equation 8.
+Analysis.
+Within a set (prompt or response), desirable behavior is (i) order preservation (no within-set crossings) and (ii) near-straight trajectories (little curvature).
+Inter-set crossings (prompt vs.
+response) are allowed—and often required—to place response tokens between prompts during infilling.
+(a) Scaled positions over time without OT.
+(b) Unscaled positions over time without OT.
+Figure 5: Ablations on number of sampling steps without OT coupling. Paths curve and cross more often, complicating learning and inference.
+(a) Scaled positions over time with OT.
+(b) Unscaled positions over time with OT.
+Figure 6: With per-set OT coupling, trajectories are straighter and order-preserving within each set.
+(a) Scaled training targets without OT: many crossings.
+(b) Scaled training targets with OT: straighter, non-crossing (within-set).
+Figure 7: Ground-truth token paths (scaled to [1,1][-1,1]) used as supervision.
+Without OT (left), many crossings indicate unstable matching; with OT (right), trajectories are nearly straight and order-preserving within each set.
+A.5 Implementation Notes
+Computing OT efficiently.
+Prompt matching is solved exactly in 1D by sorting $z_{0}^{p}$ and $z_{T}^{p}$ and pairing by rank; response matching uses order-preserving injective alignment via global-alignment DP against $z_{T}^{\mathrm{np}}$.
+Both avoid dense cost matrices and run in $O(\ell^{p}\log\ell^{p}+\ell^{r}(L-\ell^{p}))$ time and $O(L)$ memory.
+Caching & streaming. For large corpora, precompute couplings on CPU and stream mini-batches; we observed negligible overhead whether caching or computing couplings on the fly.